### v5

# Benchmarking Web API Integration Code Generation

Abstract
API integration is a cornerstone of our digital infrastructure, enabling software systems to connect and interact. However, as shown by many studies, writing or generating correct code to invoke APIs, particularly web APIs, is challenging. Although large language models (LLMs) have become popular in software development, their effectiveness in automating the generation of web API integration code remains unexplored. In order to address this, we present WAPIIBench, a dataset and evaluation pipeline designed to assess the ability of LLMs to generate web API invocation code. Our experiments with several open-source LLMs reveal that generating API invocations poses a significant challenge, resulting in hallucinated endpoints, incorrect argument usage, and other errors. None of the evaluated open-source models was able to solve more than 40% of the tasks.
I Introduction
Web APIs provide services and functionality through a standardized, HTTP-based interface. Developers leverage them to rapidly create applications (powered by so-called API integration code) that seamlessly integrate with a wide range of online services. The market surrounding API integrations is experiencing explosive growth [1, 2].
However, writing correct API integration code is a tedious and challenging task. Developers need to understand and correctly utilize a significant amount of information to be able to write an API invocation in compliance with its specification [3, 4]. Moreover, the task of ensuring correct API usage is further complicated by the fact that APIs evolve and change over time.
Large language models (LLMs) have great potential to boost productivity in software development thanks to their ability to assist developers, e.g., by automatically generating program code from natural language descriptions [5]. However, LLMs may also hallucinate and make mistakes, raising concerns about correctness [6, 7], security [8, 9], and general quality [10] of LLM-generated code.
Research gap
While the capabilities of LLMs have been investigated for many different software engineering tasks, including library (i.e., non-web) API invocations [11], their aptitude for web API integrations appears to be rather unexplored (cf. also related work in Section IV). The question is: Given the pivotal role of web APIs, can we unlock the potential of LLMs for streamlining and accelerating the development processes of software that involves web API integration? As a first step towards answering this question, we need to understand how well existing LLMs support that task. To this end, we formulate the following research question:
RQ: How correct is LLM-generated web API invocation code, and what types of errors are commonly present in such code?
Methodology
Flow diagram visualizing the evaluation pipeline described in the caption.
Figure 1: Benchmark design for evaluating the capabilities of LLMs in generating web API invocation code. 1) Based on an API $a$ and its specifications $s$, an API invocation tasks $t$ and corresponding correct request configurations $c$ are created. 2) For each $t$, the LLM under evaluation generates an API invocation ii. 3) $i$ is executed in a controlled environment, yielding a request configuration $c^{\prime}$. 4) $c^{\prime}$ is compared to $c$ and validated against $s$ to obtain various metrics, shown in Table II.
We investigate the stated research question through the following task: Given a short program context and a code comment that describes a desired operation on a given web API in natural language, the LLM has to complete the code by implementing the operation in compliance with the API’s specification.
To arrive at an answer, we create WAPIIBench, a benchmark for web API integration tasks. Specifically, we perform the following steps. First, we create a dataset of web API invocation tasks and expected outcomes. Using a powerful state-of-the-art LLM, we generate the initial version of the dataset and subject it to a thorough manual review and correction process. Second, we implement an evaluation pipeline that allows us to safely execute web API invocation code and evaluate the resulting requests to obtain fine-grained correctness metrics. For this purpose, we develop a novel evaluation methodology. Third, we use our dataset and evaluation pipeline to evaluate a comprehensive set of LLMs, including models from the StarCoder [12, 13], DeepSeek-Coder [14, 15], Qwen-Coder [16], (Code) Llama [17, 18], and GPT families.
Scope
Listing 1: General structure of an API invocation using Axios and JavaScript.
xxxxxx
const axios = require(’axios’);
axios.<method>(’http://server.com/path/to/endpoint/arg1’, {
arg2: ’request body parameters’
headers: {
arg3: ’header parameters’
params: {
arg4: ’query parameters’
Since web APIs can be used in many ways, we need to be more specific and focus our research specifically on the following setup: We target real-world web APIs that adhere to OpenAPI [^1] https://www.openapis.org/, the most widely used specification standard for web APIs [19]. We generate code in JavaScript, one of the most popular programming languages [20], and use Axios [^2] https://axios-http.com/, the leading JavaScript library for invoking web APIs [21], which allows explicit configuration of all components of an HTTP requests.
The general structure of the API invocations we consider is shown in Listing 1. An invocation comprises an HTTP method (GET, POST, etc.), a URL with server address and (potentially parameterized) path to the endpoint, a request body, a request header, and query parameters. Such web API invocations have some unique challenges compared to function calls in a local codebase: 1. Operations are identified by the combination of a method and a long URL string, not just a simple function name; 2. there is not just one but multiple argument lists and the number of available arguments can be much higher while many arguments have complex, nested data types; and 3. web API endpoints are documented externally, limiting accessibility for LLMs, and the used specification format is much more complex than, e.g., that of documentation comments used for local functions.
Evaluation results
Our study reveals that, without any in-context information about the APIs, most LLMs demonstrate some understanding of the assigned tasks, show familiarity with the general structure of invocations in JavaScript/Axios, and appear to have memorized portions of the APIs’ specifications. However, even the best open-source models generate correct invocations only 40% of the time. In particular, they frequently hallucinate endpoint URLs (up to 39%) and parameter names (up to 31%).
Contributions
In summary, by introducing WAPIIBench, we make the following contributions:
A first-of-its-kind dataset of web API invocation tasks
across four real-world APIs with nearly 400 pairs of tasks and expected outcomes, which can be used for execution-based correctness analysis. → Section II-A
An open-source evaluation pipeline
for automatically and safely evaluating the performance of code generation models on web API invocation tasks using fine-grained metrics. → Section II-B – Section II-D
Novel empirical insights
on the (in)ability of LLMs to generate correct API invocation code. → Section III
WAPIIBench is available on GitHub [^3] https://github.com/stg-tud/WAPIIBench. In addition, we provide all model-generated codes in our artifact [^4] https://doi.org/10.5281/zenodo.17607587. Additional details and comprehensive results are provided in an appendix at the end of this paper.
II Benchmark Design
WAPIIBench’s evaluation pipeline, depicted in Figure 1, consists of four stages: 1. dataset creation, 2. code generation, 3. code execution, and 4. correctness analysis. They are detailed in the following subsections.
II-A Dataset Creation
We conceive of our dataset as a collection of triples $(a,t,c)$, where $a$ is the name of the targeted API, $t$ is a task, given as an unambiguous natural language description of a request to $a$, and $c$ is a so-called request configuration, which captures all properties an API request must have in order to solve tt—it is the key to measuring functional correctness. Specifically, $c$ contains the expected URL, HTTP method, and arguments in the request body, header, and query locations. In this paper, we use the term endpoint to refer to the combination of URL and (HTTP) method.
To our knowledge, no dataset with the described structure exists. To create a new one, two general approaches exist: sourcing data from the real world or synthesizing it. We first explored the former option by searching public GitHub repositories for suitable JavaScript/Axios code [^5] Using BigQuery (https://cloud.google.com/bigquery/) for this purpose.. We dropped this option and resorted to the latter for two reasons. First, we found that real-world code rarely includes sufficient information (e.g., comments or in-code documentation) to serve as an unambiguous task description $t$. Second, the resulting dataset would have been incomplete, not covering all endpoints of a given API.
Instead, we opted for a synthetic approach, using an LLM $\operatorname{Gen}$ and API specifications $s$ to generate task–configuration pairs for our dataset: $\operatorname{Gen}(a,s)=(t,c)$. To ensure our dataset represents real-world scenarios, we chose APIs of popular services across different domains—Asana, Google Calendar, Google Sheets, and Slack—which are specified following OpenAPI, the most commonly used industry standard [19]. As $\operatorname{Gen}$, we used Gemini 1.5 Pro [^6] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ for its very large context window (2M tokens), given that OpenAPI specifications can be very long (e.g., Slack’s contains over 20,000 lines of text). Together with the full specification, we provided the model with very detailed (but mostly non-API-specific) instructions on how to construct $t$ and $c$ in a zero-shot prompt to avoid pitfalls of few-shot prompting [22] and increase generalizability. The prompt is provided in the appendix. In total, we synthesized 395 samples—one for each endpoint of the selected APIs. A detailed breakdown is shown in Table I.
To ensure the validity of the AI-generated data, we curated all samples with automated and manual checks. The automated checks caught any inconsistencies between the generated configurations $c$ and the specifications $s$. The 9 samples that failed these checks were manually inspected and corrected. Afterward, we manually checked all 395 samples to catch remaining issues that cannot be checked automatically, especially underspecification and ambiguities in the tasks $t$, in order to ensure that there is one unique $c$ for each $t$. Concretely, we checked 1. that $t$ is well-formulated and solvable; 2. that $c$ actually solves $t$; and 3. that every argument in $t$ is reflected in $c$ and vice versa. 58 samples had one or multiple issues and were fixed by adjusting $t$ and/or $c$.
Listing 2 shows a simplified example from our dataset, with the API $a$ being Google Calendar. The task $t$ is given as a comment, followed by the implementation the model has to generate. Listing 3 shows the corresponding configuration $c$, which contains key–value pairs specifying the correct HTTP method, URL, and parameters—in this case, there are only parameters in the request body (“data”).
TABLE I: Number of Endpoints per HTTP Method for Each API (Each Endpoint Corresponds to One Sample in the Dataset)
API Total POST GET PUT
DELETE
PATCH
Asana 167 61 79 14 13 0
Google Calendar 37 14 11 4 4 4
Google Sheets 17 12 4 1 0 0
Slack Web 174 95 79 0 0 0
Total 395 182 173 19 17 4
II-B Code Generation
We take each pair of API $a$ and task $t$ from our dataset and let the model under evaluation $\operatorname{M}$ generate an API invocation $i$: $\operatorname{M}(a,t)=i$. The prompt instructs $M$ to solve $t$ by completing the starter code that is given below it. We specify $a$ in order to disambiguate $t$ as otherwise a common task like “create a new user …” could refer to many different APIs. The prompt is provided in the appendix. We also include some clarifications in the prompt, meant to avoid instances of misalignment and recurring error patterns we observed in preliminary experiments. The starter code always includes $t$ as a line comment (e.g., // Create a new secondary calendar …), besides the required imports.
We conduct our experiments with two different starter code variants—we call them setups—that serve complementary purposes. Full completion includes the beginning of API invocation (axios.). By fixing the code up to this position, we ensure that an Axios call is actually generated and can be evaluated. In this setup, we can assess the ability of the model to identify the correct endpoint for $t$ and to solve the task as a whole. Argument completion additionally includes the correct method and URL (e.g., axios.post(’https://www.googleapis.com/calendar/v3/calendars’,). In this setup, we can evaluate whether the model is able to use a given endpoint in compliance with $a$. Also, this setup resembles real-world scenarios inside an IDE where the user has already implemented a part of the solution and wants it to be completed by the AI.
Note that the code generation task described here is considerably different from the dataset generation task from Section II-A. In the latter, Gemini had access to the full API specification and very detailed instructions, making this a task about processing in-context information. The models under evaluation, on the other hand, have to rely solely on memorized knowledge about the APIs, as we want to find out how well they can recall these details from their training data. This is a relevant concern as context length is limited, and it is often not desirable to fill the context with too much information. For instance, Slack’s specification has around 126K tokens, exceeding the context length of most models we evaluated.
Listing 2: Simplified task from our dataset with correct implementation.
xxxxxx
// Create a new secondary calendar named "Example Calendar" with time zone "America/Los_Angeles".
const axios = require(’axios’);
axios.post(
’https://www.googleapis.com/calendar/v3/calendars’, {
summary: ’Example Calendar’,
timeZone: ’America/Los_Angeles’
Listing 3: Request configuration corresponding to the task in Listing 2.
xxxxxx
"method": "post",
"url":
"https://www.googleapis.com/calendar/v3/calendars",
"data": {
"summary": "Example Calendar",
"timeZone": "America/Los_Angeles"
II-C Code Execution
Our approach to evaluating LLMs’ capabilities in generating API invocation code is based on functional correctness rather than syntactic similarity to a reference implementation, as the latter is prone to misjudgments [23]. To measure functional correctness, we need to execute the generated code and assess its outcome, analogous to unit testing [24].
When executing $i$, we need to address challenges unique to our setting. First, direct execution of generated code is unsafe as it could send arbitrary requests to external servers. Second, we need to capture and evaluate the configuration $c^{\prime}$ of the API request implemented by $i$, which is a side effect and not the return value of $i$. Another common challenge when evaluating code generation is excess code [25], i.e., code produced when the model terminates too late.
To address these challenges, we implement a controlled execution environment $\operatorname{Mock}$, which intercepts requests before transmission and then serializes and returns their configurations: $\operatorname{Mock}(i)=c^{\prime}$. We address the excess code issue in $\operatorname{Mock}$ by heuristically truncating the generated code after the API invocation, preventing the execution of unnecessary, incomplete, broken, or harmful code.
II-D Correctness Analysis
TABLE II: Evaluation Metrics for Correctness and Specification-Compliance of API Invocations
Metric Description
Correct implementations Generated executable code matches the ground-truth configuration
Illegal implementations Generated code contains at least one violation of the API specification
Correct URLs Generated URL matches the ground-truth URL
Illegal URLs Generated URL is not defined in the API specification
Correct methods Generated HTTP method matches the ground-truth HTTP method
Illegal methods Generated HTTP method is not defined for the generated URL in the API specification
Mean argument precision Probability that the generated arguments are correct
Mean argument recall Probability that the correct arguments are generated
Illegal arguments Generated arguments are not permitted for the generated API endpoint
Mean argument value conditional accuracy Probability that an argument value is correct if the argument name is correct
We compare the configuration $c^{\prime}$, captured in the previous step, with the ground-truth $c$ from our dataset: $c^{\prime}\overset{?}{=}c$. However, as web API requests contain multiple components—URL, method, and parameters in different locations (cf. Listing 1)—only looking for exact matches is insufficient. Instead, we compare the configurations in an element-wise manner as well as in aggregated forms to facilitate a fine-grained evaluation. Additionally, we validate $c^{\prime}$ against the API specification $s$ to distinguish between endpoints and arguments that are undesired but specification-compliant and ones that are actually illegal. The key metrics we calculate are explained in Table II and additional ones are provided in the appendix. Note that some metrics apply only to one of the setups (full/argument completion), as mentioned in Section II-B.
III Evaluation
Using our dataset and evaluation pipeline, we study the performance of 21 state-of-the-art open-source LLMs on web API invocation code generation. We primarily evaluated models that have undergone fine-tuning on code data—CodeT5+ [26] (16B), StarCoder [12] (15.5B), StarCoder2 [13] (3B, 7B, 15B), DeepSeek-Coder [14] (1.3B, 6.7B, 7B, 33B), DeepSeek-Coder-V2 [15] (16B), Qwen2.5-Coder [16] (0.5B, 1.5B, 3B, 7B, 14B, 32B), and Code Llama [17] (7B, 13B, 70B)—and additionally Llama 3.1 [18] (8B, 70B). These models were selected due to their frequent appearance in coding leaderboards [^7] E.g., https://evalplus.github.io/leaderboard.html. To obtain an upper bound on performance, we also evaluated OpenAI’s commercial GPT-4o [^8] https://openai.com/index/hello-gpt-4o/ and GPT-4o mini [^9] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ models. For a lack of space and for more visual clarity, we show only a subset of these models in our plots and tables, but our following analysis considers all of them. Comprehensive result tables are provided in the appendix.
Throughout this paper, we use the notation (t) to indicate that a metric is calculated based on the total amount of samples in the dataset and (e) to indicate that it is calculated based on only the samples that resulted in executable code [^10] To be exact, the calculation for (t) and (e) values is identical except for a filtering step to remove non-executable samples before calculating the latter.. We do this because our evaluation method can only assess the correctness of individual request components if the generated code can be executed, and would otherwise classify them as incorrect by default. The distinction between (t) and (e) allows us to differentiate between a model’s ability to generate complete and executable code (t), and a model’s ability to generate code that complies with the specification and solves the given task, provided that it is executable (e). For illustration why this distinction is necessary, consider a model $\operatorname{M}$ that predicts the correct endpoint 100% of the time but introduces syntax errors in the request arguments 50% of the time. When looking at (t) metrics only, we would wrongly conclude that $\operatorname{M}$ predicts endpoints correctly only 50% of the time.
III-A Results
[…]
Figure 2: Performance comparison between selected models for full completion (top) and argument completion (bottom).
Figure 2 shows key performance metrics (explained in Table II) for one representative [^11] With a size as close as possible to 15B for a fair comparison. of each model family plus the best-performing open-source model for full and argument completion, respectively. The corresponding numbers are provided in Tables III and IV. We used greedy decoding for all experiments.
Full completion
Across all evaluated models, the overall achieved correctness—correct implementations (t)—ranges from 0% for the Qwen2.5-Coder and Llama 3.1 models up to 60% for GPT-4o, which also dominates on all other metrics. The best open-source model is Code Llama (70B) at 30% followed by StarCoder2 at 26%. In all cases with 0% correctness, (almost) none of the generated implementations are executable—since also all other metrics for these models become zero, we omit reporting their numbers in the following. For the remaining models, the executability is 94% and higher.
Most open-source models correctly generate over 80% of HTTP methods (e). The rate of correct URLs (e) varies strongly between 32% and 72%, but 14% to 39% of URLs are illegal (e). All models have higher precision (e) than recall (e) for arguments, ranging from 51% to 75% and from 33% to 69%, respectively. If an argument name was predicted correctly, the argument value was correct 53%–84% of the time (e). For all of these metrics, Code Llama (70B) is the strongest open-source model, even surpassing GPT-4o mini. DeepSeek-Coder (7B) and CodeT5+ (16B) perform the worst across individual metrics.
Argument completion
Compared to full completion, almost all performance metrics are better for argument completion. The rate of correct implementations (t) ranges from 8% for DeepSeek-Coder (7B) to 40% for Code Llama (70B), which is again the strongest open-source model. The best performing model overall is again GPT-4o at 77% correctness. The models that struggled to generate any executable code before now achieve similar executability rates as the other models. Qwen2.5-Coder (14B) and (32B) are now even among the better-performing models.
Overall, for all metrics except correct implementations (t), we observe a smaller variance between the open-source models and an even clearer performance gap between them and GPT-4o. Since in the argument completion setup, the correct endpoint is provided to the models, we can now evaluate their ability to generate argument lists without violating the API specification. We observe that 6% to 31% of generated arguments are illegal (e).
III-B Discussion
In summary, all tested open-source LLMs struggle with generating correct API invocations. While they often generate partially correct implementations—indicating that the models have had training exposure to API specifications or examples from which they memorized usage patterns of the respective APIs—the models fail to combine individual patterns to fully correct API invocations. The full completion setup shows that the models often do not find the correct endpoint to use. But even when the correct endpoint is provided in the argument completion setup, the models still often fail to stay consistent with the given endpoint and its defined parameters.
The models perform comparatively well in predicting the HTTP method, which is expected since there is only a small set of available methods, with GET and POST being by far the most common ones (cf. Table I). In contrast, URLs and arguments are challenging to predict due to their unrestricted nature. The argument precision may be inflated to some extent by the Authorization header argument, which is required by most endpoints and therefore easy to predict. The relatively high argument value conditional accuracy indicates that identifying the correct argument is harder for the models than assigning the correct value to it once it is found.
An analysis of the generated implementations shows that those models that got nearly zero implementations correct in the full completion setup failed to understand the task given to them, resulting in incomplete or syntactically broken and thus non-executable code. Specifically, Qwen2.5-Coder mostly refused to continue the provided starter code, while Llama 3.1 skipped over the method part of the API invocation. It may be possible to solve this issue through prompt engineering, but we preferred to use a single prompt consistently for all models.
An additional finding is that larger models are not always better. In multiple model families (DeepSeek-Coder, Qwen2.5-Coder, Code Llama), we observe that the medium-sized variant performs worse than both the smaller and larger variants. Lastly, breaking down the evaluation results by API (cf. appendix), we see that each model is particularly good or bad at different APIs, which likely results from the prevalence of the respective APIs in the model’s training data. Overall, most models perform best on the Google Calendar API.
RQ: How correct is LLM-generated API invocation code, and what types of errors are commonly present in such code?
Answer: While LLMs are able to generate partially correct API invocations, the overall result is incorrect most of the time. The types of errors commonly present are manifold and range from selecting the wrong endpoint to leaving out required arguments or hallucinating entirely illegal arguments.
TABLE III: Evaluation Results for Full Completion
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Code Llama (13B) Code Llama (70B) Llama 3.1 (8B) GPT-4o
Correct implementations (t) 000.06 000.12 000.26 000.18 000.21 000.00 000.08 000.30 000.00 000.60
Correct URLs (e) 000.32 000.54 000.60 000.60 000.53 000.00 000.58 000.72 000.00 000.78
Illegal URLs (e) 000.39 000.18 000.22 000.20 000.21 000.00 000.24 000.15 000.00 000.09
Correct methods (e) 000.85 000.88 000.87 000.88 000.84 000.00 000.84 000.89 000.00 000.91
Illegal methods (e) 000.13 000.09 000.11 000.11 000.14 000.00 000.14 000.10 000.00 000.07
Mean argument precision (e) 000.52 000.65 000.68 000.65 000.67 000.00 000.72 000.75 000.00 000.89
Mean argument recall (e) 000.43 000.54 000.63 000.56 000.57 000.00 000.65 000.69 000.00 000.87
Mean argument value conditional accuracy (e) 000.53 000.68 000.79 000.72 000.76 000.00 000.57 000.84 000.00 000.95
TABLE IV: Evaluation Results for Argument Completion
CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Code Llama (13B) Code Llama (70B) Llama 3.1 (8B) GPT-4o
Correct implementations (t) 000.12 000.27 000.25 000.26 000.35 000.37 000.12 000.40 000.05 000.77
Illegal implementations (e) 000.63 000.27 000.36 000.33 000.32 000.35 000.30 000.23 000.31 000.09
Mean argument precision (e) 000.58 000.81 000.71 000.74 000.77 000.74 000.79 000.83 000.65 000.94
Mean argument recall (e) 000.53 000.74 000.65 000.67 000.71 000.72 000.74 000.79 000.43 000.94
Illegal arguments (e) 000.25 000.13 000.14 000.15 000.13 000.20 000.13 000.09 000.16 000.05
Mean argument value conditional accuracy (e) 000.71 000.83 000.83 000.83 000.86 000.78 000.65 000.89 000.67 000.95
III-C Limitations and Threats to Validity
1) Real-world API integration code is diverse and complex. To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section II-A. For instance, there is limited program context before the API invocation. Additionally, we evaluate only the correctness of the outgoing request, not how the incoming response is handled. These limitations do, however, not interfere with our investigation of LLM’s ability to recall information about correct API usage—if this information is memorized, they should be able to recall it in our evaluation as well as real-world settings.
2) The quality of the dataset and evaluation depends on the quality of the API specifications. We encountered and manually fixed several errors in the real-world specifications. Moreover, the specifications sometimes contain usage constraints that are only explained in a parameter’s free-form textual description and thus not captured by our automated correctness analysis.
3) Despite thorough vetting, our dataset might contain faulty samples that could introduce noise into the evaluation results. We also noticed that the synthetic tasks produced by Gemini 1.5 Pro tend to use optional parameters rather sparingly and often use placeholder values instead of realistic example values, limiting the transferability of findings from our benchmark to real-world API invocation tasks. Moreover, the results for some models may not generalize to other APIs, given the limited number of APIs included in our dataset. As models can be very sensitive to prompt variations, the performance of each model can almost certainly be optimized by customized prompt engineering.
IV Related Work
Differentiation of API Invocation Settings
Since APIs are such a universal concept, it is important to distinguish between different settings in which APIs are being used and understand their different characteristics when discussing related work. We divide APIs roughly into the following categories: 1. general web APIs(this work), 2. domain-specific web APIs, 3. SDK-wrapped web APIs, 4. local function APIs, and 5. tool APIs for AI agents.
1) Our work targets general web API invocations, which involve explicitly configuring all components of an HTTP request, i.e., an endpoint URL, HTTP method, and multiple request argument lists with oftentimes complex nested data types. To our knowledge, we are the first to explore and to propose a method for evaluating the capabilities of LLMs in generating code with such specific and challenging characteristics.
2) Su et al. [27, 28] and Hosseini et al. [29] investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^12] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData). Their work has a narrow focus on data-centric applications (mostly within the Microsoft ecosystem) and an SQL-like syntax. In contrast, our benchmark targets OpenAPI, a very general and flexible API standard, making the task more open-ended and challenging. Also, instead of ad-hoc API queries by a user, we strive for production-ready API integration code implemented in JavaScript.
3) Works like Dialog2API [30] and CloudAPIBench [31] focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries. Both approaches do not generalize due to SDK heterogeneity. In contrast, our benchmark relies on a unified RESTful interface for any web API. This eliminates SDK dependency, enabling direct assessment of models’ core API comprehension rather than SDK-specific knowledge. By standardizing interactions across diverse APIs, we remove evaluation biases and isolate model capabilities from implementation quirks of individual SDKs.
4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., [32, 33, 34]. Invoking a function requires specifying the function name and one argument list. In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section I (Scope).
5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment [35]. These tools range from simple utilities like calculators [36] to web services that enable AI agents to perform tasks such as weather queries or restaurant booking [37, 38, 39]. At first sight, this body of work seems related to research on automatic generation of integration code, as tools typically expose a function-like interface. However, both input and output modalities differ between these settings, as well as their non-functional requirements, so benchmarks are not directly transferable to other kinds of API invocations.
API Invocation Datasets
Since functional abstractions are essential in programming, virtually every code dataset includes (function) API invocations. Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs [27, 29], SDK-based API invocations [30, 31], and tool-using AI agents [40, 37, 38, 39, 41]. We are not aware of datasets specifically about generating API integration code to invoke REST APIs like we do.
Evaluation Methods for Web API Invocations
The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness [23], or using functional testing [24], which is not directly transferable to web API invocations, as discussed in Section II-C. This is why we developed a novel approach described in Section II. Evaluation methods in related works on API invocations mostly rely on variations of the aforementioned standard approaches. Some works use humans [42] or LLMs [39] as judges, and others execute the API invocation and verify the state of the environment [30, 41]. We are not aware of an evaluation method able to provide results as fine-grained as ours.
Studies on Memorization and Hallucination in Code LLMs
LLMs display a remarkable capacity to memorize information from their training data—which can be beneficial when it comes to generating factual responses, but also a problem when it comes to leaking private information from the training data. Studies have shown that LLMs memorize, among other things, method definitions and method calls [43], but also that they struggle to suggest the correct APIs [11]. These findings are in line with our observation that LLMs seem to know many endpoint URLs and argument names, but are often unable to assemble them to a specification-compliant API invocation that solves the task in question.
If LLMs are unable to recall memorized facts, they tend to hallucinate information that is not grounded in reality. This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments [44, 6, 7]. Our study confirms that these findings also hold for web API endpoints and their arguments, where a high amount of hallucinations can be observed.
Symbolic Approaches to Detecting API Misuse
Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues [3, 4] and real-world API misuses [45, 46, 47]. Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., [45, 48, 47]. They are, however, primarily designed to aid humans, not LLMs. Using static analyses for evaluation purposes is generally conceivable, but they can only detect whether an API invocation is legal or illegal, not whether it also solves the given task.
V Conclusion and Future Work
In this paper, we examined the challenges faced by large language models when generating reliable web API invocation code. We introduced WAPIIBench, a novel dataset and evaluation pipeline to systematically assess LLM performance in this domain. Our experiments with open-source LLMs revealed significant limitations: Only 30% (full completion) and 40% (argument completion) of the produced samples were completely right. Endpoints and arguments are often hallucinated due to insufficient recall of knowledge on API usage.
Overall, our work underscores the need for integrating quality assurance techniques into AI-driven code generation workflows to ensure correctness and reliability. Our benchmark lays the foundation for future work to investigate the impact of different approaches—such as retrieval-augmented generation [49], fine-tuning, reasoning, feedback loops, and constrained decoding [50, 51]—on the correctness and specification-compliance of generated web API invocation code. While our evaluation pipeline is specialized on JavaScript and Axios, our dataset and methodology is generic and can be transferred to other languages and libraries.
Acknowledgments
This work was funded by the Hessian Ministry of Higher Education, Research, Science and the Arts within the cluster project The Third Wave of Artificial Intelligence (3AI), by the National Research Center for Applied Cybersecurity ATHENE within the project Foundational Models for Secure Software Development, and by the LOEWE initiative (Hesse, Germany) [LOEWE/4a//519/05/00.002(0013)/95].

### v6

@@
-Benchmarking Web API Integration Code Generation
-Abstract
+Evaluating and Mitigating Errors in LLM-Generated Web API Integrations
+Abstract.
 API integration is a cornerstone of our digital infrastructure, enabling software systems to connect and interact.
@@
 None of the evaluated open-source models was able to solve more than 40% of the tasks.
-I Introduction
+Motivated by those findings, we explore the potential of constrained decoding for generating API invocations.
+To this end, we propose an automatic translation from API specifications to constraints.
+Our approach prevents violations of API usage rules and significantly increases the overall correctness of the generated code, on average by 90% and 135%, depending on the provided starter code.
+artificial intelligence, software engineering, large language models, code generation, web APIs, benchmarks, constrained decoding
+copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Automatic programmingccs: Computing methodologies Neural networksccs: Information systems RESTful web services
+1. Introduction
 Web APIs provide services and functionality through a standardized, HTTP-based interface.
 Developers leverage them to rapidly create applications (powered by so-called API integration code) that seamlessly integrate with a wide range of online services.
-The market surrounding API integrations is experiencing explosive growth [1, 2].
+The market surrounding API integrations is experiencing explosive growth (MarketsandMarkets, 2024; Postman, 2024).
 However, writing correct API integration code is a tedious and challenging task.
-Developers need to understand and correctly utilize a significant amount of information to be able to write an API invocation in compliance with its specification [3, 4].
+Developers need to understand and correctly utilize a significant amount of information to be able to write an API invocation in compliance with its specification (Robillard, 2009; Rauf et al., 2019).
 Moreover, the task of ensuring correct API usage is further complicated by the fact that APIs evolve and change over time.
-Large language models (LLMs) have great potential to boost productivity in software development thanks to their ability to assist developers, e.g., by automatically generating program code from natural language descriptions [5].
-However, LLMs may also hallucinate and make mistakes, raising concerns about correctness [6, 7], security [8, 9], and general quality [10] of LLM-generated code.
+Large language models (LLMs) have great potential to boost productivity in software development thanks to their ability to assist developers, e.g., by automatically generating program code from natural language descriptions (Cui et al., 2025).
+However, LLMs may also hallucinate and make mistakes, raising concerns about correctness (Dou et al., 2024; Tambon et al., 2025), security (Pearce et al., 2022; Perry et al., 2023), and general quality (GitClear, 2025) of LLM-generated code.
 Research gap
-While the capabilities of LLMs have been investigated for many different software engineering tasks, including library (i.e., non-web) API invocations [11], their aptitude for web API integrations appears to be rather unexplored (cf.
-also related work in Section IV).
+While the capabilities of LLMs have been investigated for many different software engineering tasks, including library (i.e., non-web) API invocations (Zhuo et al., 2023), their aptitude for web API integrations appears to be rather unexplored (cf.
+also related work in Section 7).
 The question is: Given the pivotal role of web APIs, can we unlock the potential of LLMs for streamlining and accelerating the development processes of software that involves web API integration?
@@
 To this end, we formulate the following research question:
-RQ: How correct is LLM-generated web API invocation code, and what types of errors are commonly present in such code?
-Methodology
-Flow diagram visualizing the evaluation pipeline described in the caption.
-Figure 1: Benchmark design for evaluating the capabilities of LLMs in generating web API invocation code.
-1) Based on an API $a$ and its specifications $s$, an API invocation tasks $t$ and corresponding correct request configurations $c$ are created.
-2) For each $t$, the LLM under evaluation generates an API invocation ii.
-3) $i$ is executed in a controlled environment, yielding a request configuration $c^{\prime}$.
-4) $c^{\prime}$ is compared to $c$ and validated against $s$ to obtain various metrics, shown in Table II.
+RQ1: How correct is LLM-generated web API invocation code, and what types of errors are prevalent in such code?
+Methodology for RQ1
 We investigate the stated research question through the following task: Given a short program context and a code comment that describes a desired operation on a given web API in natural language, the LLM has to complete the code by implementing the operation in compliance with the API’s specification.
@@
 For this purpose, we develop a novel evaluation methodology.
-Third, we use our dataset and evaluation pipeline to evaluate a comprehensive set of LLMs, including models from the StarCoder [12, 13], DeepSeek-Coder [14, 15], Qwen-Coder [16], (Code) Llama [17, 18], and GPT families.
+Third, we use our dataset and evaluation pipeline to evaluate a comprehensive set of LLMs, including models from the StarCoder (Li et al., 2023a; Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024b; DeepSeek-AI et al., 2024), Qwen-Coder (Hui et al., 2024), (Code) Llama (Rozière et al., 2023; Dubey et al., 2024), and GPT families.
 Scope
-Listing 1: General structure of an API invocation using Axios and JavaScript.
-xxxxxx
+Listing 1: General structure of an API invocation using Axios and JavaScript
 const axios = require(’axios’);
-axios.<method>(’http://server.com/path/to/endpoint/arg1’, {
+axios.<method>(’https://server.com/path/to/endpoint/arg1’, {
 arg2: ’request body parameters’
@@
 arg4: ’query parameters’
-Since web APIs can be used in many ways, we need to be more specific and focus our research specifically on the following setup: We target real-world web APIs that adhere to OpenAPI [^1] https://www.openapis.org/, the most widely used specification standard for web APIs [19].
-We generate code in JavaScript, one of the most popular programming languages [20], and use Axios [^2] https://axios-http.com/, the leading JavaScript library for invoking web APIs [21], which allows explicit configuration of all components of an HTTP requests.
+Since web APIs can be used in many ways, we need to be more specific and focus our research specifically on the following setup: We target real-world web APIs that adhere to OpenAPI [^1] https://www.openapis.org/, the most widely used specification standard for web APIs (SmartBear, 2020).
+We generate code in JavaScript, one of the most popular programming languages (Statista, 2024), and use Axios [^2] https://axios-http.com/, the leading JavaScript library for invoking web APIs (Greif and Burel, 2022), which allows explicit configuration of all components of an HTTP requests.
 The general structure of the API invocations we consider is shown in Listing 1.
 An invocation comprises an HTTP method (GET, POST, etc.), a URL with server address and (potentially parameterized) path to the endpoint, a request body, a request header, and query parameters.
-Such web API invocations have some unique challenges compared to function calls in a local codebase: 1.
-Operations are identified by the combination of a method and a long URL string, not just a simple function name; 2.
-there is not just one but multiple argument lists and the number of available arguments can be much higher while many arguments have complex, nested data types; and 3.
-web API endpoints are documented externally, limiting accessibility for LLMs, and the used specification format is much more complex than, e.g., that of documentation comments used for local functions.
-Evaluation results
-Our study reveals that, without any in-context information about the APIs, most LLMs demonstrate some understanding of the assigned tasks, show familiarity with the general structure of invocations in JavaScript/Axios, and appear to have memorized portions of the APIs’ specifications.
-However, even the best open-source models generate correct invocations only 40% of the time.
-In particular, they frequently hallucinate endpoint URLs (up to 39%) and parameter names (up to 31%).
+Such web API invocations have some unique challenges compared to function calls in a local codebase: (1) Operations are identified by the combination of a method and a long URL string, not just a simple function name; (2) there is not just one but multiple argument lists and the number of available arguments can be much higher while many arguments have complex, nested data types; and (3) web API endpoints are documented externally, limiting accessibility for LLMs, and the used specification format is much more complex than, e.g., that of documentation comments used for local functions.
+Evaluation results for RQ1
+Our study reveals that while most LLMs demonstrate some understanding of the assigned tasks, show familiarity with the general structure of invocations in JavaScript/Axios, and appear to have memorized portions of the APIs’ specifications, they frequently generate incorrect invocations.
+In particular, they often hallucinate endpoint URLs and parameter names.
+Based on our empirical findings, the central question shifts from whether LLMs can streamline and accelerate software development involving web API integration to how their performance on web API invocation tasks can be improved.
+Two established techniques have been extensively investigated for improving the performance of LLMs in new domains: retrieval-augmented generation (RAG) (Lewis et al., 2020), which inserts relevant information into prompts, and fine-tuning, which adapts models through domain-specific training.
+While prior research (see Section 7) indicates both methods could prove effective in our setting, each approach has notable limitations.
+RAG risks prompt bloat (Gan and Sun, 2025), which consumes limited context window capacity, obscures critical information (Liu et al., 2024a), and increases computational costs.
+Fine-tuning requires substantial computational resources and risks obsolescence when API specifications evolve.
+Importantly, neither method guarantees correctness or code quality in generated outputs.
+Given the limitations of RAG and fine-tuning approaches, this paper investigates constrained decoding (Hokamp and Liu, 2017; Deutsch et al., 2019) as an alternative method to improve reliability in LLM-generated web API integration code.
+Constrained decoding (also known as structured or guided generation) dynamically filters the token predictions of an LLM during generation to enforce compliance with predefined structural or semantic constraints.
+Crucially, this technique operates without requiring model modifications or prompt adjustments.
+To our knowledge, constrained decoding remains unexplored for web API invocations that explicitly configure all components of an HTTP request.
+To systematically study the effectiveness of constrained decoding for improving the performance of LLMs on web API invocation tasks, we formulate and answer our second research question:
+RQ2: How does constrained decoding influence the correctness of LLM-generated web API invocation code?
+Methodology for RQ2
+To answer RQ2, we perform two steps.
+First, we design a constrained decoding approach suitable for the specific use case of web API invocation tasks.
+Second, we use WAPIIBench to compare the performance of constrained and unconstrained LLMs.
+The main challenge we have to address when applying constrained decoding in our setting is constraint creation.
+Since every web API requires different constraints, we implement an automatic translation from OpenAPI specifications to constraints that are represented as regular expressions.
+When generating the constraints, the special characteristics and complexities of web API invocations need to be taken into account.
+Besides respecting the general JavaScript and the specific Axios syntax (cf.
+Listing 1), our constraints need to handle interdependencies between allowed URLs, HTTP methods, and different kinds of parameters.
+Moreover, the number, ordering, and nesting of parameters can be arbitrary, while some parameters are required and some are optional, making advanced constraint formulations necessary.
+Evaluation results for RQ2
+The results of applying our evaluation methodology to study constrained models confirm that decoding with our constraints enforces compliance with an API’s specification when generating an invocation, reliably reducing the amount of illegal URLs, HTTP methods, and arguments to zero.
+Moreover, across models, constrained decoding significantly improves the overall correctness of the generated code, at relative gains of 90% and 142% on average depending on the provided starter code.
 Contributions
-In summary, by introducing WAPIIBench, we make the following contributions:
-A first-of-its-kind dataset of web API invocation tasks
-across four real-world APIs with nearly 400 pairs of tasks and expected outcomes, which can be used for execution-based correctness analysis. → Section II-A
-An open-source evaluation pipeline
-for automatically and safely evaluating the performance of code generation models on web API invocation tasks using fine-grained metrics. → Section II-B – Section II-D
-Novel empirical insights
-on the (in)ability of LLMs to generate correct API invocation code. → Section III
-WAPIIBench is available on GitHub [^3] https://github.com/stg-tud/WAPIIBench.
-In addition, we provide all model-generated codes in our artifact [^4] https://doi.org/10.5281/zenodo.17607587.
-Additional details and comprehensive results are provided in an appendix at the end of this paper.
-II Benchmark Design
-WAPIIBench’s evaluation pipeline, depicted in Figure 1, consists of four stages: 1.
-dataset creation, 2.
-code generation, 3.
-code execution, and 4.
-correctness analysis.
+In summary, by introducing WAPIIBench and exploring constrained decoding for web API invocations, we make the following contributions:
+A first-of-its-kind dataset of web API invocation tasks:
+across four real-world APIs with nearly 400 pairs of tasks and expected outcomes, which can be used for execution-based correctness analysis. → Section 2.1
+An open-source evaluation pipeline:
+for automatically and safely evaluating the performance of code generation models on web API invocation tasks using fine-grained metrics. → Section 2.2 – Section 2.4
+A constraint generator:
+that translates OpenAPI specifications to regex-based constraints, which ensure that generated code is specification-complaint. → Section 4
+Novel empirical insights:
+on the (in)ability of LLMs to generate correct API invocation code without and with constrained decoding. → Section 3 & Section 5
+WAPIIBench is available on GitHub [^3] https://github.com/stg-tud/WAPIIBench; Note that the code also includes work-in-progress implementations for future work, which can be ignored..
+In addition, we provide all model-generated codes in our artifact [^4] https://doi.org/10.5281/zenodo.13758414.
+Additional details and comprehensive results are provided in the appendix at the end of this paper.
+2. Benchmark Design
+Refer to caption Flow diagram visualizing the evaluation pipeline described in the caption.
+Figure 1.
+Benchmark design for evaluating the capabilities of LLMs in generating web API invocation code.
+(1) Based on an API $a$ and its specifications $s$, an API invocation tasks $t$ and corresponding correct request configurations $c$ are created.
+(2) For each $t$, the LLM under evaluation generates an API invocation ii.
+(3) $i$ is executed in a controlled environment, yielding a request configuration $c^{\prime}$.
+(4) $c^{\prime}$ is compared to $c$ and validated against $s$ to obtain various metrics, shown in Table 2.
+WAPIIBench’s evaluation pipeline, depicted in Figure 1, consists of four stages: (1) dataset creation, (2) code generation, (3) code execution, and (4) correctness analysis.
 They are detailed in the following subsections.
-II-A Dataset Creation
+2.1. Dataset Creation
 We conceive of our dataset as a collection of triples $(a,t,c)$, where $a$ is the name of the targeted API, $t$ is a task, given as an unambiguous natural language description of a request to $a$, and $c$ is a so-called request configuration, which captures all properties an API request must have in order to solve tt—it is the key to measuring functional correctness.
@@
 Instead, we opted for a synthetic approach, using an LLM $\operatorname{Gen}$ and API specifications $s$ to generate task–configuration pairs for our dataset: $\operatorname{Gen}(a,s)=(t,c)$.
-To ensure our dataset represents real-world scenarios, we chose APIs of popular services across different domains—Asana, Google Calendar, Google Sheets, and Slack—which are specified following OpenAPI, the most commonly used industry standard [19].
+To ensure our dataset represents real-world scenarios, we chose APIs of popular services across different domains—Asana, Google Calendar, Google Sheets, and Slack—which are specified following OpenAPI, the most commonly used industry standard (SmartBear, 2020).
 As $\operatorname{Gen}$, we used Gemini 1.5 Pro [^6] https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ for its very large context window (2M tokens), given that OpenAPI specifications can be very long (e.g., Slack’s contains over 20,000 lines of text).
-Together with the full specification, we provided the model with very detailed (but mostly non-API-specific) instructions on how to construct $t$ and $c$ in a zero-shot prompt to avoid pitfalls of few-shot prompting [22] and increase generalizability.
+Together with the full specification, we provided the model with very detailed (but mostly non-API-specific) instructions on how to construct $t$ and $c$ in a zero-shot prompt to avoid pitfalls of few-shot prompting (Reynolds and McDonell, 2021) and increase generalizability.
 The prompt is provided in the appendix.
 In total, we synthesized 395 samples—one for each endpoint of the selected APIs.
-A detailed breakdown is shown in Table I.
+A detailed breakdown is shown in Table 1.
 To ensure the validity of the AI-generated data, we curated all samples with automated and manual checks.
@@
 Afterward, we manually checked all 395 samples to catch remaining issues that cannot be checked automatically, especially underspecification and ambiguities in the tasks $t$, in order to ensure that there is one unique $c$ for each $t$.
-Concretely, we checked 1.
-that $t$ is well-formulated and solvable; 2.
-that $c$ actually solves $t$; and 3.
-that every argument in $t$ is reflected in $c$ and vice versa.
+Concretely, we checked (1) that $t$ is well-formulated and solvable; (2) that $c$ actually solves $t$; and (3) that every argument in $t$ is reflected in $c$ and vice versa.
 58 samples had one or multiple issues and were fixed by adjusting $t$ and/or $c$.
@@
 Listing 3 shows the corresponding configuration $c$, which contains key–value pairs specifying the correct HTTP method, URL, and parameters—in this case, there are only parameters in the request body (“data”).
-TABLE I: Number of Endpoints per HTTP Method for Each API (Each Endpoint Corresponds to One Sample in the Dataset)
-API Total POST GET PUT
-DELETE
-PATCH
+Table 1. Number of endpoints per HTTP method for each API (each endpoint corresponds to one sample in the dataset)
+API Total POST GET PUT DELETE PATCH
 Asana 167 61 79 14 13 0
@@
 Total 395 182 173 19 17 4
-II-B Code Generation
+Listing 2: Simplified task from our dataset with correct implementation
+// Create a new secondary calendar named "Example Calendar" with time zone "America/Los_Angeles".
+const axios = require(’axios’);
+axios.post(’https://www.googleapis.com/calendar/v3/calendars’, {
+summary: ’Example Calendar’,
+timeZone: ’America/Los_Angeles’
+Listing 3: Request configuration corresponding to the task in Listing 2
+"method": "post",
+"url": "https://www.googleapis.com/calendar/v3/calendars",
+"data": {
+"summary": "Example Calendar",
+"timeZone": "America/Los_Angeles"
+2.2. Code Generation
 We take each pair of API $a$ and task $t$ from our dataset and let the model under evaluation $\operatorname{M}$ generate an API invocation $i$: $\operatorname{M}(a,t)=i$.
@@
 Also, this setup resembles real-world scenarios inside an IDE where the user has already implemented a part of the solution and wants it to be completed by the AI.
-Note that the code generation task described here is considerably different from the dataset generation task from Section II-A.
+Note that the code generation task described here is considerably different from the dataset generation task from Section 2.1.
 In the latter, Gemini had access to the full API specification and very detailed instructions, making this a task about processing in-context information.
@@
 For instance, Slack’s specification has around 126K tokens, exceeding the context length of most models we evaluated.
-Listing 2: Simplified task from our dataset with correct implementation.
-xxxxxx
-// Create a new secondary calendar named "Example Calendar" with time zone "America/Los_Angeles".
-const axios = require(’axios’);
-axios.post(
-’https://www.googleapis.com/calendar/v3/calendars’, {
-summary: ’Example Calendar’,
-timeZone: ’America/Los_Angeles’
-Listing 3: Request configuration corresponding to the task in Listing 2.
-xxxxxx
-"method": "post",
-"url":
-"https://www.googleapis.com/calendar/v3/calendars",
-"data": {
-"summary": "Example Calendar",
-"timeZone": "America/Los_Angeles"
-II-C Code Execution
-Our approach to evaluating LLMs’ capabilities in generating API invocation code is based on functional correctness rather than syntactic similarity to a reference implementation, as the latter is prone to misjudgments [23].
-To measure functional correctness, we need to execute the generated code and assess its outcome, analogous to unit testing [24].
+2.3. Code Execution
+Our approach to evaluating LLMs’ capabilities in generating API invocation code is based on functional correctness rather than syntactic similarity to a reference implementation, as the latter is prone to misjudgments (Evtikhiev et al., 2023).
+To measure functional correctness, we need to execute the generated code and assess its outcome, analogous to unit testing (Chen et al., 2021).
 When executing $i$, we need to address challenges unique to our setting.
@@
 Second, we need to capture and evaluate the configuration $c^{\prime}$ of the API request implemented by $i$, which is a side effect and not the return value of $i$.
-Another common challenge when evaluating code generation is excess code [25], i.e., code produced when the model terminates too late.
+Another common challenge when evaluating code generation is excess code (Guo et al., 2024a), i.e., code produced when the model terminates too late.
 To address these challenges, we implement a controlled execution environment $\operatorname{Mock}$, which intercepts requests before transmission and then serializes and returns their configurations: $\operatorname{Mock}(i)=c^{\prime}$.
 We address the excess code issue in $\operatorname{Mock}$ by heuristically truncating the generated code after the API invocation, preventing the execution of unnecessary, incomplete, broken, or harmful code.
-II-D Correctness Analysis
-TABLE II: Evaluation Metrics for Correctness and Specification-Compliance of API Invocations
+2.4. Correctness Analysis
+Table 2. Evaluation Metrics for Correctness and Specification-Compliance of API Invocations
 Metric Description
@@
 Additionally, we validate $c^{\prime}$ against the API specification $s$ to distinguish between endpoints and arguments that are undesired but specification-compliant and ones that are actually illegal.
-The key metrics we calculate are explained in Table II and additional ones are provided in the appendix.
-Note that some metrics apply only to one of the setups (full/argument completion), as mentioned in Section II-B.
-III Evaluation
+The key metrics we calculate are explained in Table 2 and additional ones are provided in the appendix.
+Note that some metrics apply only to one of the setups (full/argument completion), as mentioned in Section 2.2.
+3. Evaluation of Unconstrained LLMs
 Using our dataset and evaluation pipeline, we study the performance of 21 state-of-the-art open-source LLMs on web API invocation code generation.
-We primarily evaluated models that have undergone fine-tuning on code data—CodeT5+ [26] (16B), StarCoder [12] (15.5B), StarCoder2 [13] (3B, 7B, 15B), DeepSeek-Coder [14] (1.3B, 6.7B, 7B, 33B), DeepSeek-Coder-V2 [15] (16B), Qwen2.5-Coder [16] (0.5B, 1.5B, 3B, 7B, 14B, 32B), and Code Llama [17] (7B, 13B, 70B)—and additionally Llama 3.1 [18] (8B, 70B).
+We primarily evaluated models that have undergone fine-tuning on code data—CodeT5+ (Wang et al., 2023) (16B), StarCoder (Li et al., 2023a) (15.5B), StarCoder2 (Lozhkov et al., 2024) (3B, 7B, 15B), DeepSeek-Coder (Guo et al., 2024b) (1.3B, 6.7B, 7B, 33B), DeepSeek-Coder-V2 (DeepSeek-AI et al., 2024) (16B), Qwen2.5-Coder (Hui et al., 2024) (0.5B, 1.5B, 3B, 7B, 14B, 32B), and Code Llama (Rozière et al., 2023) (7B, 13B, 70B)—and additionally Llama 3.1 (Dubey et al., 2024) (8B, 70B).
 These models were selected due to their frequent appearance in coding leaderboards [^7] E.g., https://evalplus.github.io/leaderboard.html.
@@
 When looking at (t) metrics only, we would wrongly conclude that $\operatorname{M}$ predicts endpoints correctly only 50% of the time.
-III-A Results
+3.1. Results
 […]
-[…]
-Figure 2: Performance comparison between selected models for full completion (top) and argument completion (bottom).
-Figure 2 shows key performance metrics (explained in Table II) for one representative [^11] With a size as close as possible to 15B for a fair comparison.
+Figure 2. Performance comparison between selected models for full completion (left) and argument completion (right) with unconstrained decoding
+Figure 4 shows key performance metrics (explained in Table 2) for one representative [^11] With a size as close as possible to 15B for a fair comparison.
 of each model family plus the best-performing open-source model for full and argument completion, respectively.
-The corresponding numbers are provided in Tables III and IV.
+The corresponding numbers are provided in Tables 3 and 4.
 We used greedy decoding for all experiments.
@@
 We observe that 6% to 31% of generated arguments are illegal (e).
-III-B Discussion
+3.2. Discussion
 In summary, all tested open-source LLMs struggle with generating correct API invocations.
@@
 But even when the correct endpoint is provided in the argument completion setup, the models still often fail to stay consistent with the given endpoint and its defined parameters.
+Table 3. Evaluation results for full completion with unconstrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B) GPT-4o
+Correct implementations (t) 000.06 000.12 000.26 000.17 000.21 000.00 000.00 000.08 000.30 000.60
+Correct URLs (e) 000.32 000.54 000.60 000.60 000.53 000.00 000.00 000.58 000.72 000.78
+Illegal URLs (e) 000.39 000.18 000.22 000.20 000.21 000.00 000.00 000.24 000.15 000.09
+Correct methods (e) 000.85 000.88 000.87 000.88 000.84 000.00 000.00 000.84 000.89 000.91
+Illegal methods (e) 000.13 000.09 000.11 000.11 000.14 000.00 000.00 000.14 000.10 000.07
+Mean argument precision (e) 000.53 000.67 000.68 000.66 000.70 000.00 000.00 000.73 000.75 000.89
+Mean argument recall (e) 000.43 000.54 000.63 000.56 000.57 000.00 000.00 000.65 000.69 000.87
+Mean arg. val. conditional accuracy (e)
+000.72 000.84 000.92 000.88 000.91 000.00 000.00 000.63 000.91 000.96
+Table 4. Evaluation results for argument completion with unconstrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B) GPT-4o
+Correct implementations (t) 000.12 000.27 000.25 000.26 000.35 000.37 000.05 000.12 000.40 000.77
+Illegal implementations (e) 000.63 000.27 000.36 000.33 000.32 000.35 000.31 000.30 000.23 000.09
+Mean argument precision (e) 000.59 000.81 000.72 000.74 000.80 000.74 000.73 000.79 000.83 000.94
+Mean argument recall (e) 000.53 000.74 000.65 000.67 000.71 000.72 000.43 000.74 000.79 000.94
+Illegal arguments (e) 000.25 000.13 000.14 000.15 000.13 000.20 000.16 000.13 000.09 000.05
+Mean arg. val. conditional accuracy (e)
+000.86 000.85 000.90 000.91 000.93 000.89 000.89 000.70 000.92 000.97
 The models perform comparatively well in predicting the HTTP method, which is expected since there is only a small set of available methods, with GET and POST being by far the most common ones (cf.
-Table I).
+Table 1).
 In contrast, URLs and arguments are challenging to predict due to their unrestricted nature.
@@
 Overall, most models perform best on the Google Calendar API.
-RQ: How correct is LLM-generated API invocation code, and what types of errors are commonly present in such code?
+RQ1: How correct is LLM-generated web API invocation code, and what types of errors are prevalent in such code?
 Answer: While LLMs are able to generate partially correct API invocations, the overall result is incorrect most of the time.
 The types of errors commonly present are manifold and range from selecting the wrong endpoint to leaving out required arguments or hallucinating entirely illegal arguments.
-TABLE III: Evaluation Results for Full Completion
-CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Code Llama (13B) Code Llama (70B) Llama 3.1 (8B) GPT-4o
-Correct implementations (t) 000.06 000.12 000.26 000.18 000.21 000.00 000.08 000.30 000.00 000.60
-Correct URLs (e) 000.32 000.54 000.60 000.60 000.53 000.00 000.58 000.72 000.00 000.78
-Illegal URLs (e) 000.39 000.18 000.22 000.20 000.21 000.00 000.24 000.15 000.00 000.09
-Correct methods (e) 000.85 000.88 000.87 000.88 000.84 000.00 000.84 000.89 000.00 000.91
-Illegal methods (e) 000.13 000.09 000.11 000.11 000.14 000.00 000.14 000.10 000.00 000.07
-Mean argument precision (e) 000.52 000.65 000.68 000.65 000.67 000.00 000.72 000.75 000.00 000.89
-Mean argument recall (e) 000.43 000.54 000.63 000.56 000.57 000.00 000.65 000.69 000.00 000.87
-Mean argument value conditional accuracy (e) 000.53 000.68 000.79 000.72 000.76 000.00 000.57 000.84 000.00 000.95
-TABLE IV: Evaluation Results for Argument Completion
-CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Code Llama (13B) Code Llama (70B) Llama 3.1 (8B) GPT-4o
-Correct implementations (t) 000.12 000.27 000.25 000.26 000.35 000.37 000.12 000.40 000.05 000.77
-Illegal implementations (e) 000.63 000.27 000.36 000.33 000.32 000.35 000.30 000.23 000.31 000.09
-Mean argument precision (e) 000.58 000.81 000.71 000.74 000.77 000.74 000.79 000.83 000.65 000.94
-Mean argument recall (e) 000.53 000.74 000.65 000.67 000.71 000.72 000.74 000.79 000.43 000.94
-Illegal arguments (e) 000.25 000.13 000.14 000.15 000.13 000.20 000.13 000.09 000.16 000.05
-Mean argument value conditional accuracy (e) 000.71 000.83 000.83 000.83 000.86 000.78 000.65 000.89 000.67 000.95
-III-C Limitations and Threats to Validity
+4. Translation of API Specifications to Constraints
+Refer to caption Flow diagram visualizing the constrained decoding framework described in the caption.
+Figure 3.
+Constrained decoding framework for generating web API invocations.
+In regular autoregressive generation, a decoder takes the token sequence $s_{1},\dots,s_{i}$ and predicts the probabilities $p_{j}$ for each possible next token $t_{j}$, based on which $s_{i+1}$ is selected.
+Constrained decoding augments this loop with a constraint check that sets the probability of tokens that fail the check to zero.
+We derive our constraints from OpenAPI specifications, and they ensure that generated API invocations comply with the specification.
+Constrained decoding (Hokamp and Liu, 2017; Deutsch et al., 2019) is a general technique that modifies an LLM’s next-token predictions based on predefined constraints to prevent the generation of tokens that would lead to undesirable output sequences.
+As this technique appears to be underexplored for web API invocation code while also showing promise in related settings (cf.
+Section 7), we want to study its efficacy for web API invocation and its impact on the issues we uncovered in Section 3.
+We implemented a simple, model-agnostic constrained decoding framework for our investigations, which is depicted in Figure 3.
+The constraints are automatically derived from OpenAPI specifications and represented as regular expressions (regex).
+They are designed to ensure that generated API invocations are always compliant with the API’s specification.
+Note that our contribution here lies in this transformation of specifications into constraints, and not the implementation of the constrained decoding engine.
+While we developed a custom implementation to maintain full control over the system and to add some debugging features, migrating to an off-the-shelf engine (cf.
+Section 7) would be feasible thanks to the engine-agnostic nature of our constraints; such a migration could enhance performance and robustness.
+A main challenge when translating OpenAPI specifications to regular expressions in a generalizable and scalable way is the sheer size and complexity of real-world APIs, which is also reflected in the respective specification files.
+Each API can have dozens of endpoints (in the APIs we worked with, up to 174), and each endpoint can have dozens of parameters, distributed on different locations (body, header, path, query).
+Parameters can be optional or required, and they can be passed in arbitrary order.
+Additionally, many parameters have complex data types (objects and arrays) with a strictly defined nesting structure.
+There are also interdependencies between different parts of an API invocation, e.g., only for certain HTTP methods, parameters can be placed in the request body.
+In summary, some elements of an API invocations must be followed strictly, while other elements are flexible.
+Additional challenges arise at the syntax-level of the language the API invocation is implemented in.
+Our regexes should accept valid JavaScript while not limiting variations permitted by the language, such as formatting or the use of different quotation marks.
+This is important as different LLMs show different stylistic preferences—forcing them away from their preferred style may degrade performance.
+There are also a lot of edge cases that need to be handled carefully by the constraints, e.g., an escaped quotation mark must not terminate a string literal, and many other seemingly minor details—our experiments showed that models will find and exploit such loopholes.
+We address the challenges by structuring constraints hierarchically from the HTTP method and the URL to the different parameters.
+Once the method is matched, the search space is reduced to URLs for which an endpoint with the respective method is defined.
+Next, once the URL is matched, the search space is reduced to the parameters of the identified endpoint.
+Finally, once the parameter name is matched, the search space is reduced to values with the right data type.
+Within the parameter lists of a given endpoint, we use capture groups [^12] https://docs.python.org/3/library/re.html to keep track of parameters that have been defined.
+Conditional patternsfootnotemark: are then used to check the existence of a specific group and thus guard against the double definition of a parameter.
+Moreover, conditionals are used in conjunction with negative lookahead assertionsfootnotemark: to make it impossible to close a parameter list before all required parameters have been defined.
+While we could represent all our constraints as one monolithic regex, we opted for a more modular and flexible approach [^13] They could still be joined if needed for compatibility with other frameworks..
+We organize constraints as a set of so-called generation rules.
+Each generation rule has a regex-encoded start condition, stop condition, and body.
+Once the start condition matches the code generated so far, we constrain the generation according to the body until the stop condition matches.
+There is one specialized generation rule for each HTTP method, which is triggered when the corresponding Axios invocation is generated (e.g., axios.post().
+As multiple generation rules may be active at the same time (a token is only admissible if no rule rejects it), we can leverage overlapping constraints.
+We do so via a special generation rule that becomes active before all other rules and guides the generation towards the start condition of the other rules, ensuring one of them is going to be activated (thus avoiding using methods not supported by the given API).
+Moreover, we let generation rule bodies overlap with their stop conditions to ensure they properly terminate.
+This approach can also be generalized to many other applications of constrained decoding to orchestrate multiple constraints in a flexible and efficient manner.
+5. Evaluation of Constrained LLMs
+We put our constraints into action with all open-source models [^14] As they are closed-source, we cannot apply constrained decoding to the GPT models.
+under test and reuse WAPIIBench’s dataset and evaluation pipeline from Section 2.
+Again, we focus the analysis in this section on a subset of models, but provide comprehensive results in the appendix.
+5.1. Results
+[…]
+Figure 4. Performance comparison between selected models for full completion (left) and argument completion (right) with constrained decoding
+Refer to caption Stacked bar chart visualizing the performance difference between unconstrained and constrained decoding.
+Figure 5. Relative increase in correct codes through constrained decoding for full (left) and argument completion (right)
+Figure 4 shows key performance metrics for a selection of models, now equipped with constrained decoding.
+Additionally, Figure 5 shows the rate of correct codes (t) with and without constrained decoding.
+Full completion
+Constrained decoding increases the correctness across metrics and models.
+Code Llama (70B) remains the strongest open-source model at 46% correct codes (t).
+The relative correctness gains through constrained decoding vary strongly between the models but are generally high, at an average of +90% and up to +200% for Code Llama (13B).
+Qwen2.5-Coder and Llama 3.1, which were previously unable to generate any executable code, now achieve executability rates similar to the other models, while there are slightly more non-executable codes overall.
+Notably, the number of illegal URLs and HTTP methods is reduced to zero.
+Argument completion
+In contrast to the unconstrained baseline, where the performance difference across metrics between the full and argument completion setups was not very pronounced, we observe a much more distinct jump in performance when employing constrained decoding as well as a reduced variance between the models.
+The relative gains in correct codes (t) average at +135% and go up to +879% for Llama 3.1 (8B).
+Notably, constrained decoding reduces the number of illegal arguments to zero for all models.
+Table 5. Evaluation results for full completion with constrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B)
+Correct implementations (t) 000.13 000.15 000.40 000.29 000.27 000.07 000.19 000.25 000.46
+Correct URLs (e) 000.36 000.56 000.67 000.66 000.56 000.26 000.40 000.60 000.75
+Illegal URLs (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
+Correct methods (e) 000.86 000.89 000.93 000.93 000.92 000.40 000.70 000.92 000.93
+Illegal methods (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
+Mean argument precision (e) 000.64 000.76 000.78 000.79 000.77 000.58 000.65 000.76 000.86
+Mean argument recall (e) 000.66 000.70 000.79 000.77 000.73 000.64 000.66 000.78 000.84
+Mean arg. val. conditional accuracy (e) 000.82 000.71 000.94 000.94 000.94 000.74 000.88 000.81 000.96
+Table 6. Evaluation results for argument completion with constrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (15B) DeepSeek-Coder (6.7B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (14B) Llama 3.1 (8B) Code Llama (13B) Code Llama (70B)
+Correct implementations (t) 000.35 000.31 000.49 000.47 000.48 000.46 000.47 000.44 000.63
+Illegal implementations (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
+Mean argument precision (e) 000.90 000.95 000.92 000.92 000.93 000.93 000.91 000.93 000.96
+Mean argument recall (e) 000.86 000.90 000.87 000.90 000.89 000.93 000.89 000.94 000.93
+Illegal arguments (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00
+Mean arg. val. conditional accuracy (e) 000.89 000.81 000.93 000.96 000.96 000.84 000.93 000.86 000.97
+5.2. Discussion
+In summary, constrained decoding significantly improves the correctness of generated API invocations by eliminating illegal (i.e., hallucinated) URLs, methods, and arguments—a crucial advancement toward reliable generation of API invocation code.
+The better efficacy for argument completion is expected, as fixing URLs and methods allows the constraints to substantially narrow down the search space.
+This suggests that constrained decoding may be especially well suited for in-IDE code completion tasks, where parts of the code are already provided by the user.
+Models of all sizes benefit from constrained decoding; it especially shapes up to be an expedient way to make mid-size LLMs competitive to larger ones or to make open-source models viable alternatives to commercial closed-source models (e.g., Code Llama (70B) with constrained decoding matches the performance of GPT-4o mini).
+However, it remains to be investigated why the quantitative benefit provided by constrained decoding is so model-dependent.
+In general, we believe that constrained decoding has the potential to be a powerful tool for reducing specific types of errors in code model outputs.
+It effectively prunes the model’s search space using background knowledge, particularly valuable for complex, requirement-heavy tasks.
+A key advantage is that constrained decoding guarantees compliance with specified constraints, whereas unconstrained models may inconsistently recall critical information from their training data.
+RQ2: How does constrained decoding influence the correctness of LLM-generated web API invocation code?
+Answer: Constrained decoding is an effective method to improve the overall correctness and reliability of generated API invocations.
+It provides relative correctness gains of +90% and +135% on average, depending on the setup.
+Crucially, it guarantees the absence of illegal URLs, HTTP methods, and arguments.
+6. Limitations and Threats to Validity
 1) Real-world API integration code is diverse and complex.
-To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section II-A.
+To be able to evaluate API invocations automatically, we had to apply some restrictions and simplifications when creating the tasks in our dataset, as described in Section 2.1.
 For instance, there is limited program context before the API invocation.
@@
 These limitations do, however, not interfere with our investigation of LLM’s ability to recall information about correct API usage—if this information is memorized, they should be able to recall it in our evaluation as well as real-world settings.
-2) The quality of the dataset and evaluation depends on the quality of the API specifications.
+2) The quality of the dataset, evaluation, and constraints depends on the quality of the API specifications.
 We encountered and manually fixed several errors in the real-world specifications.
-Moreover, the specifications sometimes contain usage constraints that are only explained in a parameter’s free-form textual description and thus not captured by our automated correctness analysis.
+Moreover, the specifications sometimes contain usage constraints that are only explained in a parameter’s free-form textual description and thus not captured by our automated correctness analysis or constraint generator.
+Future work could use an NLP-based approach to extract these kinds of constraints.
 3) Despite thorough vetting, our dataset might contain faulty samples that could introduce noise into the evaluation results.
@@
 As models can be very sensitive to prompt variations, the performance of each model can almost certainly be optimized by customized prompt engineering.
-IV Related Work
+4) Our constraints act only locally from the beginning of an API invocation to its end.
+When (instead of literals) variables are used as parameter values, we have no control over such a variable’s value or data type.
+To overcome this, we would need to extend our approach with a backtracking mechanism, which comes with its own set of issues (Jiang et al., 2025).
+5) Our implementation requires the user to provide the name and specification of the API, for which invocations are to be generated.
+The limitation could be addressed by automatically inferring the API referred to in a task description and retrieving the corresponding specification.
+However, our work targets use cases in the API integrations industry, where usually the customer knows which APIs they want to be integrated, and the developer only has to decide which endpoints to use in order to connect the APIs in the desired way.
+In this scenario, having to provide the correct API specification to our system is not a real limitation.
+Our work is not intended for highly exploratory development scenarios, but future work could fill this gap.
+6) The executability rate of the code generated by constrained models is slightly lower than that of unconstrained models, which can be attributed to misalignment between the LLMs’ subword tokens and the constraints.
+This can lead a model to generate tokens it would not have chosen by itself, potentially causing erratic behavior by distorting the model’s next token probability distribution (Beurer-Kellner et al., 2024).
+This issue can be resolved by switching to a minimally invasive constrained decoding engine (Beurer-Kellner et al., 2024; Ugare et al., 2025).
+However, sporadic cases of non-executable codes are no big concern for practical applications like IDEs, where code suggestions can be generated in the background and discarded or regenerated if they fail a syntax check.
+7. Related Work
 Differentiation of API Invocation Settings
 Since APIs are such a universal concept, it is important to distinguish between different settings in which APIs are being used and understand their different characteristics when discussing related work.
-We divide APIs roughly into the following categories: 1.
-general web APIs(this work), 2.
-domain-specific web APIs, 3.
-SDK-wrapped web APIs, 4.
-local function APIs, and 5.
-tool APIs for AI agents.
+We divide APIs roughly into the following categories: (1) general web APIs(this work), (2) domain-specific web APIs, (3) SDK-wrapped web APIs, (4) local function APIs, and (5) tool APIs for AI agents.
 1) Our work targets general web API invocations, which involve explicitly configuring all components of an HTTP request, i.e., an endpoint URL, HTTP method, and multiple request argument lists with oftentimes complex nested data types.
@@
 2) Su et al.
-[27, 28] and Hosseini et al.
-[29] investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^12] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData).
+(Su et al., 2017, 2018) and Hosseini et al.
+(Hosseini et al., 2021) investigated natural language interfaces to domain-specific web APIs following the Open Data Protocol [^15] https://docs.oasis-open.org/odata/odata/v4.0/odata-v4.0-part1-protocol.html (OData).
 Their work has a narrow focus on data-centric applications (mostly within the Microsoft ecosystem) and an SQL-like syntax.
@@
 Also, instead of ad-hoc API queries by a user, we strive for production-ready API integration code implemented in JavaScript.
-3) Works like Dialog2API [30] and CloudAPIBench [31] focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries.
+3) Works like Dialog2API (Shu et al., 2022) and CloudAPIBench (Jain et al., 2025) focused on SDK-wrapped web API invocations (e.g., for AWS), treating web APIs like local libraries.
 Both approaches do not generalize due to SDK heterogeneity.
@@
 By standardizing interactions across diverse APIs, we remove evaluation biases and isolate model capabilities from implementation quirks of individual SDKs.
-4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., [32, 33, 34].
+4) Local function APIs are the most common kind of API in typical codebases and studied in many works, e.g., (Zan et al., 2025; Eghbali and Pradel, 2024; Dutta et al., 2024).
 Invoking a function requires specifying the function name and one argument list.
-In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section I (Scope).
-5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment [35].
-These tools range from simple utilities like calculators [36] to web services that enable AI agents to perform tasks such as weather queries or restaurant booking [37, 38, 39].
+In contrast, general web API invocations are subject to more complex syntax and semantics than such local function calls, as already mentioned in Section 1 (Scope).
+5) There is a large body of work concerned with enabling LLM-based AI agents to use tools to interact with their environment (Yehudai et al., 2025).
+These tools range from simple utilities like calculators (Schick et al., 2023) to web services that enable AI agents to perform tasks such as weather queries or restaurant booking (Song et al., 2023; Li et al., 2023b; Qin et al., 2024).
 At first sight, this body of work seems related to research on automatic generation of integration code, as tools typically expose a function-like interface.
@@
 Since functional abstractions are essential in programming, virtually every code dataset includes (function) API invocations.
-Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs [27, 29], SDK-based API invocations [30, 31], and tool-using AI agents [40, 37, 38, 39, 41].
+Code datasets involving web APIs can be found in the fields of natural language interfaces to web APIs (Su et al., 2017; Hosseini et al., 2021), SDK-based API invocations (Shu et al., 2022; Jain et al., 2025), and tool-using AI agents (Patil et al., 2024; Song et al., 2023; Li et al., 2023b; Qin et al., 2024; Yan et al., 2024).
 We are not aware of datasets specifically about generating API integration code to invoke REST APIs like we do.
 Evaluation Methods for Web API Invocations
-The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness [23], or using functional testing [24], which is not directly transferable to web API invocations, as discussed in Section II-C.
-This is why we developed a novel approach described in Section II.
+The correctness of AI-generated code is traditionally measured either using exact matches or syntax-based similarity, which both correlate weakly with actual correctness (Evtikhiev et al., 2023), or using functional testing (Chen et al., 2021), which is not directly transferable to web API invocations, as discussed in Section 2.3.
+This is why we developed a novel approach described in Section 2.
 Evaluation methods in related works on API invocations mostly rely on variations of the aforementioned standard approaches.
-Some works use humans [42] or LLMs [39] as judges, and others execute the API invocation and verify the state of the environment [30, 41].
+Some works use humans (Ding et al., 2025) or LLMs (Qin et al., 2024) as judges, and others execute the API invocation and verify the state of the environment (Shu et al., 2022; Yan et al., 2024).
 We are not aware of an evaluation method able to provide results as fine-grained as ours.
@@
 LLMs display a remarkable capacity to memorize information from their training data—which can be beneficial when it comes to generating factual responses, but also a problem when it comes to leaking private information from the training data.
-Studies have shown that LLMs memorize, among other things, method definitions and method calls [43], but also that they struggle to suggest the correct APIs [11].
+Studies have shown that LLMs memorize, among other things, method definitions and method calls (Yang et al., 2024), but also that they struggle to suggest the correct APIs (Zhuo et al., 2023).
 These findings are in line with our observation that LLMs seem to know many endpoint URLs and argument names, but are often unable to assemble them to a specification-compliant API invocation that solves the task in question.
 If LLMs are unable to recall memorized facts, they tend to hallucinate information that is not grounded in reality.
-This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments [44, 6, 7].
+This is a common error of LLMs, and studies have shown that it also exists in code generation, where, among other things, models may hallucinate function names and arguments (Liu et al., 2024b; Dou et al., 2024; Tambon et al., 2025).
 Our study confirms that these findings also hold for web API endpoints and their arguments, where a high amount of hallucinations can be observed.
 Symbolic Approaches to Detecting API Misuse
-Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues [3, 4] and real-world API misuses [45, 46, 47].
-Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., [45, 48, 47].
+Correct usage of APIs is challenging also for humans, as indicated by several studies on API usability issues (Robillard, 2009; Rauf et al., 2019) and real-world API misuses (Bae et al., 2014; Li et al., 2021; Wang and Zhao, 2023).
+Several (static analysis) methods have been proposed for detecting (and mitigating) API misuses, e.g., (Bae et al., 2014; Binder et al., 2022; Wang and Zhao, 2023).
 They are, however, primarily designed to aid humans, not LLMs.
 Using static analyses for evaluation purposes is generally conceivable, but they can only detect whether an API invocation is legal or illegal, not whether it also solves the given task.
-V Conclusion and Future Work
+Constrained Decoding for Reliable Code Generation
+Constrained decoding for code generation commonly enforces syntactic validity using general-purpose constrained decoding frameworks (Willard and Louf, 2023; Beurer-Kellner et al., 2024; Ugare et al., 2025; Dong et al., 2024).
+These frameworks guide the generation based on regular expressions or context-free grammars, and many of them leverage precomputed finite-state machines to optimize performance.
+Since syntax errors make up only a small part of errors in LLM-generated code (Dou et al., 2024; Tambon et al., 2025), several works have explored constrained decoding to prevent more advanced kinds of errors.
+PICARD (Scholak et al., 2021) constrains SQL statements to use only existing tables and columns.
+Synchromesh (Poesia et al., 2022) uses hand-crafted semantic constraints to make generated DSL code obey certain constrains, e.g., related to scopes and types.
+Another application (Mündler et al., 2025) uses rules from the TypeScript type system to ensure that generated code is correctly typed.
+While our constraints do not incorporate a complete type system, we do ensure that arguments passed as literals match the type given in the OpenAPI specification.
+Several works have created constraints specifically to mitigate hallucinations related to function/method invocations, which appears to be a common issue (Dou et al., 2024; Tambon et al., 2025).
+Monitor-Guided Decoding (MGD) (Agrawal et al., 2023) and Repilot (Wei et al., 2023) use constraints provided by a Java language server to prevent various errors, including the hallucination of methods dereferenced from an object.
+MGD additionally constrains the number of method arguments.
+MARIN (Chen et al., 2025) mines dependencies in Java programs to combine constrained decoding and RAG at method call sites, preventing hallucinations of methods and their arguments.
+The method calls targeted by these three works have a much simpler structure than web API invocations targeted by our work, as already discussed at the beginning of this section.
+Moreover, language servers and dependency mining can only leverage information contained within the code base, but web APIs are specified externally.
+These approaches are thus in general unsuitable for our use case.
+MGD shares one feature with our approach in that it also supports multiple constraints (so-called monitors) with individual start and stop conditions.
+We go, however, one step further and facilitate the interplay between different constraints for smoother transitions between unconstrained and constrained decoding as described in Section 4.
+Related works from the line of research on AI agents have applied constrained decoding to ensure that tool APIs are used correctly.
+ToolDec (Zhang et al., 2023) uses constrained decoding for tools based on Python functions and REST APIs.
+However, tools are invoked via a simple JSON-like syntax.
+FANTASE (Wang et al., 2024) combines constrained decoding of tool names and arguments with RAG (Lewis et al., 2020) and a ranked beam search of candidate tool invocations.
+Invocations use Python syntax.
+Compared to these works, our work deals with structurally more complex API invocations, necessitating more sophisticated constraint formulations.
+Alternative Approaches to Support LLMs during API Invocations
+Besides constrained decoding, several other approaches have been proposed to mitigate problems LLMs face when generating API invocations.
+Most of these works focus on local libraries (primarily in Python), but some also consider web APIs, even though they are not as complex as the ones targeted by our work (cf.
+discussion at the beginning of this section).
+These approaches include iterative refinement (Eghbali and Pradel, 2024; Ding et al., 2025), retrieval-augmented generation (RAG) (Patil et al., 2024; Zan et al., 2025; Zhou et al., 2023; Jain et al., 2025), and fine-tuning (Patil et al., 2024; Zan et al., 2025; Dutta et al., 2024) to reduce hallucinations and improve correctness.
+The most important difference between these previous works and ours is that none of them—iterative refinement, RAG, and fine-tuning—can provide guarantees (additional disadvantages, such as higher computation cost, prompt bloat, and outdated training data, are discussed in Section 1).
+In contrast, our approach based on constrained decoding guarantees the absence of illegal endpoint URLs, HTTP methods, and request arguments, thereby improving reliability.
+Pointing out constrained decoding as a powerful part of the solution is an important contribution of our work.
+However, we believe that all the approaches discussed here should be seen as complementary, promising synergistic effects when combined.
+8. Conclusion and Future Work
 In this paper, we examined the challenges faced by large language models when generating reliable web API invocation code.
@@
 Endpoints and arguments are often hallucinated due to insufficient recall of knowledge on API usage.
+To address these challenges, we proposed the use of constrained decoding—a method that does not require costly retraining or additional context—for generating web API invocations.
+We developed a custom translation algorithm that derives constraints, represented as regular expressions, from OpenAPI specifications.
+This algorithm effectively manages the complexity and scale of real-world APIs, enabling the enforcement of compliance during code generation.
+Our results demonstrate that constrained decoding, when equipped with appropriate constraints, can eliminate illegal and hallucinated endpoints and arguments by enforcing compliance with API specifications, substantially improving the accuracy and reliability of LLM-generated API code.
+Furthermore, our translation algorithm is implementation-agnostic and can be easily integrated as a plugin into other coding assistant systems to automatically enforce API specifications during development.
 Overall, our work underscores the need for integrating quality assurance techniques into AI-driven code generation workflows to ensure correctness and reliability.
-Our benchmark lays the foundation for future work to investigate the impact of different approaches—such as retrieval-augmented generation [49], fine-tuning, reasoning, feedback loops, and constrained decoding [50, 51]—on the correctness and specification-compliance of generated web API invocation code.
+While our focus has been on constrained decoding, our benchmark lays the foundation for future work to investigate the impact of further approaches—such as retrieval-augmented generation, fine-tuning, reasoning, or feedback loops—on the correctness and specification-compliance of generated web API invocation code.
 While our evaluation pipeline is specialized on JavaScript and Axios, our dataset and methodology is generic and can be transferred to other languages and libraries.
-Acknowledgments
+Finally, we aim to extend constrained decoding to additional AI code generation tasks, broadening its impact and applicability.
+Acknowledgements.
 This work was funded by the Hessian Ministry of Higher Education, Research, Science and the Arts within the cluster project The Third Wave of Artificial Intelligence (3AI), by the National Research Center for Applied Cybersecurity ATHENE within the project Foundational Models for Secure Software Development, and by the LOEWE initiative (Hesse, Germany) [LOEWE/4a//519/05/00.002(0013)/95].
+Appendix A Data Availability
+WAPIIBench is available on GitHub [^16] https://github.com/stg-tud/WAPIIBench; Note that the code also includes work-in-progress implementations for future work, which can be ignored..
+In addition, we provide all model-generated codes in our artifact [^17] https://doi.org/10.5281/zenodo.13758414.
+Appendix B Background
+openapi: 3.0.0
+servers:
+- url: https://www.googleapis.com/calendar/v3
+info:
+title: Calendar API
+description: Manipulates events and other calendar data.
+paths:
+/calendars:
+post:
+description: Creates a secondary calendar.
+parameters:
+- name: prettyPrint
+description: Returns response with indentations and line breaks.
+in: query
+required: false
+schema:
+type: boolean
+requestBody:
+content:
+application/json:
+schema:
+properties:
+summary:
+description: Title of the calendar.
+type: string
+timeZone:
+description: The time zone of the calendar. (Formatted as an IANA Time Zone Database name, e.g. "Europe/Zurich".) Optional.
+type: string
+type: object
+security:
+- Oauth2:
+- https://www.googleapis.com/auth/calendar
+(a) OpenAPI specification (YAML)
+// Create a secondary calendar with summary "Example Calendar" and time zone "America/Los_Angeles". Pretty print the response.
+const axios = require(’axios’);
+axios.post(’https://www.googleapis.com/calendar/v3/calendars’, {
+summary: ’Example Calendar’,
+timeZone: ’America/Los_Angeles’,
+headers: {
+Authorization: ’Bearer <access_token>’
+params: {
+prettyPrint: true,
+}).then(response => {
+console.log(’Calendar created’, response.data);
+(b) API invocation (JavaScript)
+"headers": {
+"Accept": "application/json, text/plain, */*",
+"Content-Type": "application/json",
+"Authorization": "Bearer <access_token>"
+"params": {
+"prettyPrint": true
+"method": "post",
+"url": "https://www.googleapis.com/calendar/v3/calendars",
+"data": {
+"summary": "Example Calendar",
+"timeZone": "America/Los_Angeles"
+(c) Configuration object (JSON)
+Figure 6.
+Example of a web API and its usage.
+Top: Excerpt from the Google Calendar OpenAPI specification.
+Left: JavaScript code to send a request to this API using the Axios library.
+Right: Configuration object that describes the request sent.
+For our evaluation, we pair the task description (comment in the JavaScript code) with the configuration to create an input–output sample.
+Three code listings, exemplifying an OpenAPI specification, an API invocation, and a configuration object.
+OpenAPI provides a structured, human- and machine-readable way to document an API’s endpoints, requests, response formats, authentication methods, and other details.
+For this paper, we define the term endpoint as a unique combination of path and HTTP method.
+Specification files are written in JSON or YAML.
+Listing 6(a) shows an excerpt of the OpenAPI specification of the Google Calendar API [^18] Adapted from https://api.apis.guru/v2/specs/googleapis.com/calendar/v3/openapi.yaml; the original comprises 3190 lines..
+Besides some meta information, it contains a list of paths (e.g., /calendars) and, for each path, a list of supported HTTP methods (e.g., post).
+Each endpoint has properties for documentation purposes and a list of named parameters (prettyPrint).
+Parameters can be passed in different locations, indicated by the property in: path parameters are inserted directly into of the URL, query parameters are appended to the URL as key–value pairs, and header parameters become part of the HTTP request header.
+Additionally, methods like POST use the requestBody to transfer data to the server (e.g., summary, timeZone).
+We consider this data as another kind of parameter.
+Parameters have a schema that specifies their data type and constraints on permissible values.
+The security property determines available authentication schemes, such as OAuth 2.0.
+Listing 6(b) shows how a request to the endpoint from the specification in Listing 6(a) could be implemented in JavaScript using the Axios library.
+The method called on the axios object determines the HTTP method, and the first argument determines the server URL and path.
+This is followed by one object containing the request body and another object containing header and query parameters (headers, params).
+Query parameters are automatically serialized and appended to the URL by Axios.
+If we execute the code in Listing 6(b), a configuration object as shown in Listing 6(c) is created and a request, configured accordingly, is sent to the given URL.
+The configuration contains all arguments explicitly given in the code, as well as some implicit parameters, such as Accept (for the expected response media type) and Content-Type (for the request body media type).
+The representation of requests as configuration objects is key to our evaluation method described in the paper.
+Appendix C Constrained Decoding Details
+We implemented our constrained decoding engine as an extension of the Hugging Face Transformers library by providing a LogitsProcessor subclass that manages the whole constrained decoding process.
+The implementation is model-agnostic and compatible with all encoder-decoder and decoder-only architectures available on Hugging Face.
+Our constraints ensure for each argument that its name is defined in the specification for the location it appears in, and that the associated expression is either a valid variable name or a value that matches the specified schema and data type.
+Complex data structures such as arrays and objects are constrained deeply, i.e., constraints are constructed for their members as well, recursively.
+Moreover, we ensure that an argument is not passed more than once in the same location and that required arguments are indeed passed.
+While the general call syntax is predetermined in favor of the most common way to use Axios (even though alternative syntaxes exist), whitespace is handled flexibly by our constraints to permit different code formattings.
+When using greedy decoding, where always the most likely next token is generated, we optimize the search for allowed next tokens by checking candidate tokens in order of their probability and stop once the first matching token is found.
+Further performance optimizations are out of scope for this work, as related work has already proposed effective measures to improve it (Poesia et al., 2022; Willard and Louf, 2023; Beurer-Kellner et al., 2024; Ugare et al., 2025; Dong et al., 2024), which could be incorporated into our implementation.
+Our approach requires the user to select the API he or she wants to use and provide its specification.
+For an automated API selection, more research on processing the task description and classifying the implied API would be necessary.
+The evaluation in this work was limited to one API at a time.
+Enforcing constraints for multiple different APIs at once requires only implementing OR-semantics in addition to the current AND-semantics for generation rules.
+The generation rules we implemented are activated when axios.
+is generated and deactivated when the end of a statement (e.g., by ); or ).then() is generated.
+This makes it feasible to generate whole programs while only constraining Axios invocations.
+However, both the start and the stop conditions are heuristics and may fail, e.g., if a matching string is found in a comment or nested expression.
+Improving the detection of start and stop conditions is only possible to a certain extent due to the fundamental limitations of regular expressions (e.g., it is impossible to decide if parentheses are balanced).
+When generating whole programs, we also cannot constrain the data types and values of variables that are declared before and used within the Axios invocation, as it cannot be foreseen which variable will be used for what (and if it is relevant to the API invocation in the first place).
+More research on this problem is needed for constrained decoding approaches in general.
+Our constraints are designed to match a code snippet if it contains an Axios invocation that respects JavaScript syntax and API usage constraints from an OpenAPI specification.
+However, due to some simplifications we made to keep the complexity of the generated regexes in bounds, the opposite is not true, i.e., there are invalid Axios invocations that are matched by our constraints.
+One syntax-related example are single, double, and backtick quotation marks, for which we do not always ensure that the opening and the closing one match.
+However, we did not encounter any cases where a model made such a mistake.
+Hence, we did not consider it worthwhile to implement such constraints, which would add even more complexity to the regexes.
+Another OpenAPI-related example is the schema property of any parameter, which provides additional constraints beyond its data type, e.g., that a string value should be an email address or that an integer value should be in the range from 0 to 100.
+Although it would be feasible to respect such schema constraints (at least approximately), we did not implement them so far, since by constraining the data type, we already cover the majority of parameter-value-related mistakes.
+Finally, OpenAPI specifications sometimes contain usage constraints that are only explained in a parameter’s description property, which contains a free form textual description of an operation.
+For instance, the maxResults parameter in the Google Calendar API contains a passage “The page size can never be larger than 250 entries”, but this information is not reflected in the parameter’s schema (even though schemas support specifying maximum and minimum values for a parameter).
+More research is needed to enable extracting such constraints from descriptions and using them for constrained decoding and evaluation purposes.
+For API parameters that are nested objects or arrays, some simplifications were necessary.
+Since OpenAPI permits recursive schemas, we prevent an infinite regress while generating constraints from self-recursive schemas by limiting the recursion depth to three.
+This depth is sufficient for all samples in our dataset but could be adapted if necessary.
+For arrays, we had to set the maximum number of elements to five, due to an incompatibility with the way we prevent duplicate arguments.
+This is, however, no fundamental limitation.
+Again, the chosen maximum is sufficient for our dataset, and we argue that arrays of more than five elements should anyway not be defined inline in most cases.
+Appendix D Technologies
+We implemented WAPIIBench using the following technologies:
+Appendix E Models
+As can be seen, we use only base (i.e., non-instruction-tuned) models.
+While instruction-tuned models tend to perform better on coding tasks [^19] Cf., e.g., the EvalPlus leaderboard (https://evalplus.github.io/leaderboard.html), we considered them to be inappropriate for our setting, which is based on code completion.
+Performing code completion on instruction-tuned models leads to rather unnatural results, as the models often do not directly return the completion and instead start with some response (“Here is your completed code …”) followed by a code snipped that may contain (1) only the completion, (2) the starter code followed by the completion, or (3) an arbitrarily modified version of the starter code and corresponding completion.
+These factors make the model output very hard to parse reliably.
+Therefore, we decided to focus our evaluation on base models.
+The following hyperparameters were used:
+  • floating point precision = 16 bit
+  • # beams = 1
+  • temperature = 0.0
+Appendix F Prompts
+Listing 4: Prompt for generating the test data for a given API. Words in curly braces are placeholders.
+Consider this OpenAPI specification:
+‘‘‘yaml
+{spec}
+I want you to generate test data for this API. The data should be in JSON format and look like this:
+‘‘‘json
+"samples": [
+"task": "...",
+"config": {
+"url": "...",
+"method": "...",
+"headers": { ... },
+"params": { ... },
+"data": { ... }
+* samples is an array containing all the test cases.
+* task is a natural language description of a specific task that can be solved by sending a request to the given API.
+All information required to unambiguously identify and implement the corresponding API request must be contained in the task description.
+Therefore, state all expected argument values explicitly.
+The only exception is authentication keys and tokens, which must not be specified here (but you may specify the authentication method to be used if multiple ones are available).
+* config is an object containing the expected configuration of the described request. If a property of config is empty, it can be omitted.
+* url is the full URL of the endpoint (server URL + path) and may include path parameters but no query parameters.
+* method is the HTTP method used. It can be one of get‘, put‘, post‘, delete‘, and patch‘.
+* headers is an object containing all expected header arguments.
+Note that the property ‘"Accept": "application/json, text/plain, */*"‘ is always present and if data is sent in the request body, ‘"Content-Type": "mime/type"‘ is present as well (substitute "mime/type" for the respective media type).
+* params is an object containing all expected query arguments.
+* data is an object containing all data from the request body. It is only used for the methods put, post, delete, and patch.
+Remember that for authentication, an additional argument might be required.
+Take a look at an endpoints security property and the securitySchemes section in the specification, to find out if and how to authenticate.
+I assume you know how the usual authentication schemes apiKey‘, http‘, and oauth2 work.
+Use ‘<key>‘ as a placeholder for API keys (e.g., ‘"name": "<key>"‘) and ‘<token>‘ as a placeholder for authorization tokens (e.g., ‘"Authorization": "Bearer <token>"‘).
+{api_specific_instructions}
+Now, please give me a JSON object that matches my description and that contains diverse examples of requests to this API.
+{path_selection}
+Listing 5: Prompt for generating API invocations given a task description. Words in curly braces are placeholders.
+You are an AI programming assistant that helps users write API requests.
+You are given a comment that describes what the user wants to achieve and are supposed to implement it using the Axios library in JavaScript.
+For this, write a single call to Axios (with syntax axios.method(url[, config])‘) that does exactly what was described in the comment.
+* Make sure to include all parameters in config that are required to solve the given task but dont include any unnecessary parameters.
+* Insert all values directly into the place where they belong, rather than using intermediate variables.
+* If the API requires some form of authentication, use ‘<key>‘ as a placeholder for API keys or ‘<token>‘ as a placeholder for authorization tokens, respectively.
+* If a request body requires a media type other than text/json‘, explicitly set the Content-Type header to the respective type, and Axios will automatically serialize the request body accordingly.
+Your next task is about the {api} API. Complete the following code snippet{extra_instructions}:
+‘‘‘javascript
+// {task}
+const axios = require(’axios’);
+axios.{method}(’{url}’,
+Listing 4 shows the prompt used for creating the dataset with Gemini 1.5 pro.
+To avoid incomplete responses when generating the dataset based on OpenAPI specifications with more than 100 endpoints (Asana and Slack), we prompted it multiple times, asking for different subsets of endpoints.
+The Slack API required additional instructions to account for changes in the API that are not reflected in the OpenAPI specification [^20] https://api.slack.com/changelog/2017-10-keeping-up-with-the-jsons.
+Listing 5 shows the prompt used in our evaluation pipeline when generating API invocation code.
+In both cases, we decided against few-shot prompting to increase generalizability and to avoid associated pitfalls.
+Few-shot prompting techniques mainly help the model to understand the task it is supposed to solve and the desired response format.
+This is not the issue in the type of task we investigate.
+Rather, the main problem is that models select the wrong API endpoints and/or pass the wrong arguments to the endpoint.
+On the other hand, using few-shot prompts introduces new challenges, such as a high sensitivity to the examples provided while reducing the generalizability of the prompt and not guaranteeing inherently superior performance to that of zero-shot prompting (Reynolds and McDonell, 2021)
+Further optimization of model performance via prompt engineering would require either revealing information about the API’s endpoints and parameters—undermining the purpose of our evaluation of the models’ memorized knowledge—or tailoring prompts to each individual model (e.g., adjusting wording, formatting, etc.).
+However, this approach relies on trial-and-error and introduces high uncertainty and variance.
+Appendix G Extended Results
+Table 7. Complete evaluation metrics for correctness and specification-compliance of API invocations
+Metric Description
+Executable codes Generated code is complete and contains no syntax or runtime error
+Correct codes Generated executable code matches the ground-truth configuration
+Illegal codes Generated code contains at least one violation of the API specification
+Correct URLs Generated URL matches the ground-truth URL
+Illegal URLs Generated URL is not defined in the API specification
+Correct methods Generated HTTP method matches the ground-truth HTTP method
+Illegal methods Generated HTTP method is not defined for the generated URL in the API specification
+Correct argument names Generated arguments are correct
+Correct argument values Generated argument values are correct
+Missing arguments Expected arguments are not generated
+Unexpected arguments Generated arguments are not expected
+Unnecessary arguments Redundant arguments are generated for an API endpoint
+Illegal arguments Generated arguments are not permitted for the generated API endpoint
+Mean argument precision Probability that the generated arguments are correct
+Mean argument recall Probability that the correct arguments are generated
+Mean argument Jaccard index Overlap between generated and correct arguments
+Mean argument value conditional accuracy Probability that an argument value is correct if the argument name is correct
+Total errors Any type of error prevented execution
+Incomplete codes Generated code did not contain a complete API invocation
+Runtime errors Generated code produced an error when trying to execute it
+Timeouts Constrained decoding was canceled because finding allowed next tokens took too long
+Unsatisfiable constraints Constrained decoding was canceled because no allowed next token existed
+The full set of metrics calculated by our pipeline is explained in Table 7.
+Tables 8 to 11 provide a comprehensive summary of these metrics for all evaluated models [^21] While we also evaluated Gemini 1.5 Pro on WAPIIBench, we excluded it from all discussion—since it was the model that generated the dataset, its evaluation results might be skewed.
+(cf.
+Appendix E).
+Additionally, the results are broken down by API exemplarily for StarCoder2 (Lozhkov et al., 2024) and GPT-4o in Tables 12 to 17.
+The best values are in bold.
+Tables 18(a) to 18(d) compare the correctness rates of all models with unconstrained vs.
+constrained decoding.
+Note that some metrics can only be calculated either for full completion or for argument completion and are therefore not shown in all tables.
+Values marked with (t) are ratios relative to the 395 total samples, while values marked with (e) are relative to the subset of executable codes, which varies from experiment to experiment.
+The complete raw data these results are based on can be found in our artifact (cf.
+Appendix A).
+Table 8. Complete evaluation results for full completion with unconstrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (3B) StarCoder2 (7B) StarCoder2 (15B) DeepSeek-Coder (1.3B) DeepSeek-Coder (6.7B) DeepSeek-Coder (7B) DeepSeek-Coder (33B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (0.5B) Qwen2.5-Coder (1.5B) Qwen2.5-Coder (3B) Qwen2.5-Coder (7B) Qwen2.5-Coder (14B) Qwen2.5-Coder (32B) Llama 3.1 (8B) Llama 3.1 (70B) Code Llama (7B) Code Llama (13B) Code Llama (70B) Gemini 1.5 Pro GPT-4o mini GPT-4o Executable implementations (t) 000.95 000.95 000.97 000.95 000.99 000.96 000.96 000.57 000.98 000.95 000.01 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.99 000.89 000.99 1.00 1.00 1.00 Correct implementations (t) 000.06 000.12 000.11 000.10 000.26 000.08 000.17 000.04 000.17 000.21 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.12 000.08 000.30 000.45 000.39 000.60 Correct implementations (e) 000.06 000.13 000.12 000.11 000.27 000.08 000.18 000.06 000.17 000.21 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.12 000.09 000.31 000.45 000.39 000.60 Correct URLs (t) 000.31 000.52 000.40 000.45 000.60 000.33 000.57 000.35 000.67 000.51 000.01 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.48 000.51 000.72 000.72 000.57 000.78 Correct URLs (e) 000.32 000.54 000.41 000.48 000.60 000.35 000.60 000.62 000.68 000.53 000.67 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.48 000.58 000.72 000.72 000.57 000.78 Illegal URLs (t) 000.37 000.17 000.22 000.27 000.22 000.32 000.19 000.10 000.13 000.20 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.28 000.21 000.15 000.19 000.23 000.09 Illegal URLs (e) 000.39 000.18 000.23 000.29 000.22 000.34 000.20 000.17 000.13 000.21 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.28 000.24 000.15 000.20 000.23 000.09 Correct methods (t) 000.81 000.84 000.78 000.88 000.86 000.79 000.84 000.53 000.86 000.81 000.01 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.73 000.75 000.88 000.91 000.83 000.91 Correct methods (e) 000.85 000.88 000.81 000.93 000.87 000.83 000.88 000.93 000.87 000.84 1.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.74 000.84 000.89 000.92 000.83 000.91 Illegal methods (t) 000.12 000.09 000.12 000.05 000.11 000.12 000.10 000.03 000.10 000.13 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.22 000.13 000.10 000.07 000.14 000.07 Illegal methods (e) 000.13 000.09 000.12 000.05 000.11 000.12 000.11 000.05 000.10 000.14 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.23 000.14 000.10 000.07 000.14 000.07 Correct argument names (t) 000.39 000.50 000.49 000.41 000.63 000.45 000.52 000.25 000.58 000.52 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.37 000.56 000.68 000.79 000.76 000.86 Correct argument names (e) 000.41 000.53 000.51 000.43 000.63 000.47 000.55 000.48 000.59 000.55 000.29 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.38 000.63 000.68 000.80 000.76 000.87 Correct argument values (t) 000.29 000.43 000.40 000.33 000.58 000.39 000.47 000.17 000.53 000.48 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.33 000.39 000.63 000.76 000.73 000.83 Correct argument values (e) 000.30 000.45 000.42 000.35 000.58 000.41 000.49 000.33 000.54 000.50 000.29 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.33 000.43 000.63 000.77 000.73 000.83 Missing arguments (t) 000.61 000.50 000.51 000.59 000.37 000.55 000.48 000.75 000.42 000.48 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 000.63 000.44 000.32 000.21 000.24 000.14 Missing arguments (e) 000.59 000.47 000.49 000.57 000.37 000.53 000.45 000.52 000.41 000.45 000.71 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.62 000.37 000.32 000.20 000.24 000.13 Unexpected arguments (t) 000.30 000.25 000.34 000.31 000.25 000.32 000.25 000.11 000.24 000.22 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.17 000.22 000.20 000.18 000.17 000.11 Unexpected arguments (e) 000.31 000.26 000.35 000.32 000.25 000.33 000.26 000.20 000.25 000.23 000.12 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.17 000.24 000.20 000.18 000.17 000.11 Mean argument precision (t) 000.53 000.67 000.55 000.54 000.68 000.55 000.66 000.73 000.63 000.70 000.67 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.76 000.73 000.75 000.82 000.82 000.89 Mean argument precision (e) 000.53 000.67 000.55 000.54 000.68 000.55 000.66 000.73 000.63 000.70 000.67 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.76 000.73 000.75 000.82 000.82 000.89 Mean argument recall (t) 000.41 000.51 000.48 000.42 000.63 000.47 000.54 000.28 000.56 000.55 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.40 000.58 000.69 000.81 000.78 000.87 Mean argument recall (e) 000.43 000.54 000.50 000.44 000.63 000.49 000.56 000.49 000.57 000.57 000.33 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.41 000.65 000.69 000.82 000.78 000.87 Mean arg.
+Jaccard index (t) 000.34 000.46 000.40 000.35 000.56 000.38 000.48 000.25 000.47 000.49 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.37 000.52 000.62 000.75 000.72 000.83 Mean arg.
+Jaccard index (e) 000.36 000.48 000.41 000.37 000.56 000.40 000.50 000.44 000.48 000.52 000.33 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.37 000.58 000.63 000.75 000.72 000.83 Mean arg.
+val.
+cond.
+acc.
+(t) 000.72 000.84 000.80 000.84 000.92 000.85 000.88 000.69 000.91 000.91 1.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.87 000.63 000.91 000.97 000.96 000.96 Mean arg.
+val.
+cond.
+acc.
+(e) 000.72 000.84 000.80 000.84 000.92 000.85 000.88 000.69 000.91 000.91 1.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.87 000.63 000.91 000.97 000.96 000.96 Total errors 18.00 20.00 13.00 21.00 3.00 16.00 16.00 170.00 6.00 18.00 392.00 395.00 393.00 395.00 395.00 390.00 395.00 395.00 5.00 42.00 2.00 1.00 0.00 1.00 Incomplete implementations 14.00 3.00 11.00 5.00 2.00 12.00 3.00 2.00 6.00 9.00 29.00 35.00 12.00 2.00 160.00 0.00 4.00 0.00 4.00 5.00 1.00 1.00 0.00 0.00 Runtime errors 4.00 17.00 2.00 16.00 1.00 4.00 13.00 168.00 0.00 9.00 363.00 360.00 381.00 393.00 235.00 390.00 391.00 395.00 1.00 37.00 1.00 0.00 0.00 1.00
+Table 9. Complete evaluation results for argument completion with unconstrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (3B) StarCoder2 (7B) StarCoder2 (15B) DeepSeek-Coder (1.3B) DeepSeek-Coder (6.7B) DeepSeek-Coder (7B) DeepSeek-Coder (33B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (0.5B) Qwen2.5-Coder (1.5B) Qwen2.5-Coder (3B) Qwen2.5-Coder (7B) Qwen2.5-Coder (14B) Qwen2.5-Coder (32B) Llama 3.1 (8B) Llama 3.1 (70B) Code Llama (7B) Code Llama (13B) Code Llama (70B) Gemini 1.5 Pro GPT-4o mini GPT-4o Executable implementations (t) 000.94 000.97 000.98 000.99 000.99 000.97 000.99 000.99 000.99 000.98 000.89 000.95 000.92 000.99 000.93 000.99 000.98 000.99 000.99 000.95 1.00 1.00 000.99 1.00 Correct implementations (t) 000.12 000.27 000.21 000.19 000.25 000.21 000.26 000.08 000.25 000.35 000.08 000.03 000.16 000.28 000.37 000.38 000.05 000.29 000.23 000.12 000.40 000.61 000.63 000.77 Correct implementations (e) 000.13 000.27 000.22 000.19 000.25 000.22 000.27 000.08 000.25 000.36 000.09 000.03 000.18 000.28 000.40 000.38 000.05 000.29 000.23 000.13 000.40 000.61 000.64 000.77 Illegal implementations (t) 000.59 000.27 000.48 000.47 000.35 000.51 000.32 000.21 000.38 000.31 000.58 000.50 000.44 000.41 000.32 000.32 000.30 000.26 000.11 000.28 000.23 000.17 000.20 000.09 Illegal implementations (e) 000.63 000.27 000.49 000.47 000.36 000.52 000.33 000.22 000.39 000.32 000.65 000.53 000.48 000.41 000.35 000.32 000.31 000.26 000.11 000.30 000.23 000.17 000.20 000.09 Correct argument names (t) 000.49 000.71 000.62 000.53 000.64 000.60 000.66 000.52 000.69 000.67 000.43 000.47 000.44 000.67 000.65 000.80 000.43 000.63 000.48 000.70 000.78 000.90 000.88 000.93 Correct argument names (e) 000.52 000.73 000.63 000.54 000.65 000.62 000.67 000.53 000.69 000.69 000.49 000.50 000.48 000.67 000.70 000.81 000.44 000.64 000.48 000.73 000.79 000.90 000.89 000.93 Correct argument values (t) 000.42 000.61 000.52 000.44 000.57 000.52 000.60 000.45 000.64 000.62 000.34 000.41 000.40 000.63 000.58 000.73 000.38 000.58 000.42 000.51 000.72 000.85 000.84 000.90 Correct argument values (e) 000.45 000.62 000.53 000.44 000.58 000.54 000.61 000.46 000.65 000.63 000.40 000.43 000.43 000.63 000.63 000.74 000.39 000.58 000.42 000.53 000.72 000.85 000.84 000.90 Missing arguments (t) 000.51 000.29 000.38 000.47 000.36 000.40 000.34 000.48 000.31 000.33 000.57 000.53 000.56 000.33 000.35 000.20 000.57 000.37 000.52 000.30 000.22 000.10 000.12 000.07 Missing arguments (e) 000.48 000.27 000.37 000.46 000.35 000.38 000.33 000.47 000.31 000.31 000.51 000.50 000.52 000.33 000.30 000.19 000.56 000.36 000.52 000.27 000.21 000.10 000.11 000.07 Unnecessary arguments (t) 000.03 000.05 000.08 000.07 000.07 000.05 000.07 000.06 000.07 000.03 000.03 000.03 000.03 000.06 000.03 000.07 000.03 000.04 000.05 000.05 000.05 000.04 000.02 000.01 Unnecessary arguments (e) 000.03 000.05 000.08 000.07 000.07 000.05 000.07 000.06 000.07 000.03 000.03 000.03 000.03 000.06 000.03 000.07 000.03 000.04 000.05 000.05 000.05 000.04 000.02 000.01 Illegal arguments (t) 000.24 000.13 000.21 000.20 000.14 000.23 000.14 000.08 000.13 000.13 000.29 000.24 000.23 000.17 000.18 000.11 000.16 000.12 000.06 000.13 000.09 000.09 000.09 000.05 Illegal arguments (e) 000.25 000.13 000.21 000.20 000.14 000.24 000.15 000.09 000.13 000.13 000.31 000.25 000.24 000.17 000.20 000.11 000.16 000.12 000.06 000.13 000.09 000.09 000.09 000.05 Mean argument precision (t) 000.59 000.81 000.62 000.62 000.72 000.65 000.74 000.81 000.72 000.80 000.54 000.64 000.64 000.72 000.74 000.78 000.73 000.81 000.86 000.79 000.83 000.88 000.90 000.94 Mean argument precision (e) 000.59 000.81 000.62 000.62 000.72 000.65 000.74 000.81 000.72 000.80 000.54 000.64 000.64 000.72 000.74 000.78 000.73 000.81 000.86 000.79 000.83 000.88 000.90 000.94 Mean argument recall (t) 000.50 000.72 000.61 000.53 000.64 000.63 000.66 000.52 000.66 000.70 000.46 000.46 000.46 000.68 000.67 000.78 000.42 000.66 000.53 000.70 000.78 000.90 000.90 000.93 Mean argument recall (e) 000.53 000.74 000.62 000.54 000.65 000.64 000.67 000.52 000.67 000.71 000.52 000.49 000.50 000.68 000.72 000.79 000.43 000.67 000.53 000.74 000.79 000.90 000.90 000.94 Mean arg.
+Jaccard index (t) 000.42 000.66 000.51 000.46 000.57 000.54 000.59 000.46 000.57 000.64 000.38 000.39 000.41 000.59 000.62 000.71 000.38 000.61 000.49 000.64 000.72 000.85 000.85 000.91 Mean arg.
+Jaccard index (e) 000.45 000.68 000.52 000.47 000.57 000.55 000.60 000.47 000.57 000.65 000.43 000.41 000.45 000.59 000.67 000.71 000.39 000.61 000.49 000.67 000.73 000.85 000.86 000.92 Mean arg.
+val.
+cond.
+acc.
+(t) 000.86 000.85 000.82 000.85 000.90 000.86 000.91 000.90 000.93 000.93 000.80 000.87 000.90 000.95 000.89 000.92 000.89 000.91 000.88 000.70 000.92 000.96 000.96 000.97 Mean arg.
+val.
+cond.
+acc.
+(e) 000.86 000.85 000.82 000.85 000.90 000.86 000.91 000.90 000.93 000.93 000.80 000.87 000.90 000.95 000.89 000.92 000.89 000.91 000.88 000.70 000.92 000.96 000.96 000.97 Total errors 23.00 11.00 9.00 4.00 3.00 11.00 4.00 5.00 3.00 6.00 45.00 19.00 32.00 2.00 28.00 2.00 9.00 5.00 3.00 19.00 1.00 1.00 2.00 1.00 Incomplete implementations 22.00 5.00 7.00 4.00 1.00 11.00 4.00 3.00 3.00 6.00 44.00 17.00 31.00 2.00 27.00 2.00 6.00 1.00 3.00 5.00 1.00 1.00 0.00 0.00 Runtime errors 1.00 6.00 2.00 0.00 2.00 0.00 0.00 2.00 0.00 0.00 1.00 2.00 1.00 0.00 1.00 0.00 3.00 4.00 0.00 14.00 0.00 0.00 2.00 1.00
+Table 10. Complete evaluation results for full completion with constrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (3B) StarCoder2 (7B) StarCoder2 (15B) DeepSeek-Coder (1.3B) DeepSeek-Coder (6.7B) DeepSeek-Coder (7B) DeepSeek-Coder (33B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (0.5B) Qwen2.5-Coder (1.5B) Qwen2.5-Coder (3B) Qwen2.5-Coder (7B) Qwen2.5-Coder (14B) Qwen2.5-Coder (32B) Llama 3.1 (8B) Llama 3.1 (70B) Code Llama (7B) Code Llama (13B) Code Llama (70B) Executable implementations (t) 000.84 000.89 000.81 000.86 000.90 000.87 000.82 000.69 000.90 000.84 000.68 000.80 000.81 000.84 000.87 000.78 000.82 000.80 000.74 000.88 000.85 Correct implementations (t) 000.13 000.15 000.18 000.18 000.40 000.18 000.29 000.08 000.39 000.27 000.04 000.02 000.07 000.30 000.07 000.36 000.19 000.31 000.23 000.25 000.46 Correct implementations (e) 000.16 000.17 000.23 000.21 000.45 000.21 000.35 000.12 000.44 000.32 000.06 000.02 000.09 000.36 000.08 000.46 000.24 000.38 000.31 000.28 000.54 Correct URLs (t) 000.30 000.50 000.41 000.52 000.61 000.40 000.54 000.45 000.61 000.47 000.15 000.18 000.23 000.46 000.23 000.51 000.32 000.42 000.42 000.53 000.64 Correct URLs (e) 000.36 000.56 000.51 000.61 000.67 000.46 000.66 000.66 000.68 000.56 000.22 000.23 000.28 000.55 000.26 000.65 000.40 000.52 000.57 000.60 000.75 Illegal URLs (t) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Illegal URLs (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Correct methods (t) 000.72 000.79 000.72 000.82 000.83 000.75 000.77 000.61 000.81 000.76 000.33 000.31 000.39 000.64 000.34 000.56 000.57 000.63 000.64 000.81 000.79 Correct methods (e) 000.86 000.89 000.89 000.94 000.93 000.87 000.93 000.88 000.90 000.92 000.49 000.39 000.48 000.76 000.40 000.71 000.70 000.79 000.87 000.92 000.93 Illegal methods (t) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Illegal methods (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Correct argument names (t) 000.51 000.60 000.55 000.58 000.68 000.60 000.61 000.40 000.70 000.58 000.39 000.44 000.41 000.60 000.53 000.64 000.48 000.55 000.50 000.65 000.69 Correct argument names (e) 000.63 000.69 000.70 000.69 000.77 000.71 000.75 000.59 000.78 000.71 000.58 000.54 000.51 000.73 000.61 000.84 000.61 000.73 000.71 000.75 000.82 Correct argument values (t) 000.42 000.45 000.51 000.45 000.64 000.48 000.57 000.28 000.67 000.55 000.21 000.28 000.29 000.54 000.41 000.60 000.43 000.53 000.44 000.55 000.67 Correct argument values (e) 000.52 000.52 000.64 000.54 000.72 000.57 000.70 000.42 000.75 000.67 000.30 000.35 000.36 000.67 000.47 000.80 000.54 000.71 000.63 000.63 000.79 Missing arguments (t) 000.49 000.40 000.45 000.42 000.32 000.40 000.39 000.60 000.30 000.42 000.61 000.56 000.59 000.40 000.47 000.36 000.52 000.45 000.50 000.35 000.31 Missing arguments (e) 000.37 000.31 000.30 000.31 000.23 000.29 000.25 000.41 000.22 000.29 000.42 000.46 000.49 000.27 000.39 000.16 000.39 000.27 000.29 000.25 000.18 Unexpected arguments (t) 000.30 000.21 000.26 000.20 000.20 000.26 000.19 000.18 000.19 000.19 000.32 000.36 000.34 000.22 000.35 000.16 000.29 000.16 000.22 000.23 000.12 Unexpected arguments (e) 000.35 000.23 000.31 000.23 000.22 000.29 000.22 000.24 000.21 000.22 000.41 000.41 000.39 000.25 000.38 000.20 000.35 000.20 000.29 000.26 000.14 Mean argument precision (t) 000.64 000.76 000.70 000.79 000.78 000.71 000.79 000.73 000.79 000.77 000.59 000.52 000.58 000.77 000.58 000.81 000.65 000.82 000.75 000.76 000.86 Mean argument precision (e) 000.64 000.76 000.70 000.79 000.78 000.71 000.79 000.73 000.79 000.77 000.59 000.52 000.58 000.77 000.58 000.81 000.65 000.82 000.75 000.76 000.86 Mean argument recall (t) 000.55 000.62 000.59 000.62 000.71 000.65 000.63 000.43 000.72 000.61 000.42 000.45 000.44 000.64 000.55 000.68 000.54 000.61 000.54 000.69 000.71 Mean argument recall (e) 000.66 000.70 000.73 000.71 000.79 000.75 000.77 000.62 000.80 000.73 000.61 000.56 000.54 000.76 000.64 000.87 000.66 000.76 000.74 000.78 000.84 Mean arg.
+Jaccard index (t) 000.43 000.55 000.48 000.55 000.63 000.54 000.56 000.37 000.64 000.54 000.31 000.31 000.34 000.57 000.42 000.61 000.44 000.56 000.47 000.60 000.67 Mean arg.
+Jaccard index (e) 000.52 000.62 000.60 000.63 000.70 000.62 000.69 000.54 000.72 000.65 000.46 000.39 000.42 000.67 000.48 000.77 000.54 000.70 000.63 000.68 000.78 Mean arg.
+val.
+cond.
+acc.
+(t) 000.82 000.71 000.91 000.77 000.94 000.76 000.94 000.70 000.97 000.94 000.52 000.57 000.69 000.90 000.74 000.94 000.88 000.96 000.89 000.81 000.96 Mean arg.
+val.
+cond.
+acc.
+(e) 000.82 000.71 000.91 000.77 000.94 000.76 000.94 000.70 000.97 000.94 000.52 000.57 000.69 000.90 000.74 000.94 000.88 000.96 000.89 000.81 000.96 Total errors 65.00 45.00 77.00 54.00 40.00 52.00 70.00 122.00 41.00 65.00 125.00 80.00 74.00 63.00 53.00 86.00 73.00 80.00 104.00 47.00 59.00 Incomplete implementations 61.00 21.00 43.00 14.00 14.00 18.00 37.00 102.00 16.00 26.00 91.00 46.00 44.00 19.00 31.00 0.00 29.00 30.00 79.00 17.00 44.00 Runtime errors 3.00 23.00 29.00 36.00 24.00 34.00 29.00 20.00 25.00 32.00 31.00 31.00 28.00 42.00 13.00 20.00 39.00 31.00 25.00 28.00 15.00 Timeouts 0.00 1.00 1.00 4.00 2.00 0.00 4.00 0.00 0.00 0.00 2.00 0.00 0.00 2.00 8.00 66.00 5.00 1.00 0.00 2.00 0.00 Unsatisfiable constraints 1.00 0.00 4.00 0.00 0.00 0.00 0.00 0.00 0.00 7.00 1.00 3.00 2.00 0.00 1.00 0.00 0.00 18.00 0.00 0.00 0.00
+Table 11. Complete evaluation results for argument completion with constrained decoding
+CodeT5+ (16B) StarCoder (15.5B) StarCoder2 (3B) StarCoder2 (7B) StarCoder2 (15B) DeepSeek-Coder (1.3B) DeepSeek-Coder (6.7B) DeepSeek-Coder (7B) DeepSeek-Coder (33B) DeepSeek-Coder-V2 (16B) Qwen2.5-Coder (0.5B) Qwen2.5-Coder (1.5B) Qwen2.5-Coder (3B) Qwen2.5-Coder (7B) Qwen2.5-Coder (14B) Qwen2.5-Coder (32B) Llama 3.1 (8B) Llama 3.1 (70B) Code Llama (7B) Code Llama (13B) Code Llama (70B) Executable implementations (t) 000.84 000.90 000.91 000.93 000.93 000.94 000.87 000.70 000.91 000.89 000.83 000.92 000.88 000.88 000.90 000.89 000.88 000.86 000.79 000.92 000.89 Correct implementations (t) 000.35 000.31 000.37 000.30 000.49 000.39 000.47 000.23 000.58 000.48 000.14 000.11 000.27 000.51 000.46 000.65 000.47 000.51 000.38 000.44 000.63 Correct implementations (e) 000.41 000.34 000.41 000.32 000.53 000.41 000.55 000.32 000.64 000.55 000.17 000.12 000.31 000.58 000.50 000.73 000.53 000.59 000.48 000.47 000.70 Illegal implementations (t) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Illegal implementations (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Correct argument names (t) 000.70 000.80 000.82 000.78 000.79 000.85 000.77 000.52 000.83 000.76 000.71 000.82 000.73 000.78 000.84 000.84 000.76 000.75 000.66 000.85 000.81 Correct argument names (e) 000.84 000.89 000.91 000.84 000.86 000.92 000.89 000.77 000.92 000.88 000.86 000.91 000.83 000.91 000.92 000.96 000.88 000.89 000.86 000.93 000.92 Correct argument values (t) 000.63 000.65 000.74 000.61 000.75 000.71 000.74 000.49 000.81 000.73 000.50 000.59 000.60 000.75 000.71 000.81 000.71 000.73 000.61 000.74 000.79 Correct argument values (e) 000.75 000.73 000.82 000.66 000.81 000.76 000.86 000.72 000.90 000.85 000.61 000.65 000.69 000.87 000.78 000.93 000.82 000.86 000.79 000.82 000.90 Missing arguments (t) 000.30 000.20 000.18 000.22 000.21 000.15 000.23 000.48 000.17 000.24 000.29 000.18 000.27 000.22 000.16 000.16 000.24 000.25 000.34 000.15 000.19 Missing arguments (e) 000.16 000.11 000.09 000.16 000.14 000.08 000.11 000.23 000.08 000.12 000.14 000.09 000.17 000.09 000.08 000.04 000.12 000.11 000.14 000.07 000.08 Unnecessary arguments (t) 000.10 000.07 000.16 000.11 000.08 000.10 000.09 000.10 000.09 000.08 000.10 000.11 000.10 000.08 000.09 000.06 000.10 000.05 000.09 000.09 000.04 Unnecessary arguments (e) 000.12 000.08 000.17 000.11 000.09 000.11 000.10 000.14 000.10 000.09 000.12 000.12 000.11 000.09 000.10 000.07 000.11 000.06 000.12 000.10 000.05 Illegal arguments (t) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Illegal arguments (e) 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 000.00 Mean argument precision (t) 000.90 000.95 000.86 000.92 000.92 000.92 000.92 000.88 000.91 000.93 000.92 000.92 000.92 000.93 000.93 000.94 000.91 000.96 000.91 000.93 000.96 Mean argument precision (e) 000.90 000.95 000.86 000.92 000.92 000.92 000.92 000.88 000.91 000.93 000.92 000.92 000.92 000.93 000.93 000.94 000.91 000.96 000.91 000.93 000.96 Mean argument recall (t) 000.72 000.82 000.83 000.80 000.81 000.87 000.78 000.55 000.83 000.79 000.72 000.84 000.74 000.80 000.84 000.86 000.79 000.77 000.69 000.86 000.83 Mean argument recall (e) 000.86 000.90 000.91 000.85 000.87 000.93 000.90 000.78 000.92 000.89 000.87 000.91 000.84 000.92 000.93 000.97 000.89 000.90 000.87 000.94 000.93 Mean arg.
+Jaccard index (t) 000.66 000.78 000.73 000.74 000.76 000.82 000.73 000.50 000.79 000.74 000.67 000.79 000.69 000.76 000.79 000.82 000.73 000.74 000.64 000.81 000.80 Mean arg.
+Jaccard index (e) 000.79 000.86 000.81 000.80 000.81 000.87 000.84 000.72 000.87 000.84 000.81 000.86 000.79 000.86 000.88 000.92 000.83 000.87 000.81 000.88 000.90 Mean arg.
+val.
+cond.
+acc.
+(t) 000.89 000.81 000.91 000.79 000.93 000.80 000.96 000.94 000.98 000.96 000.70 000.68 000.83 000.97 000.84 000.97 000.93 000.97 000.92 000.86 000.97 Mean arg.
+val.
+cond.
+acc.
+(e) 000.89 000.81 000.91 000.79 000.93 000.80 000.96 000.94 000.98 000.96 000.70 000.68 000.83 000.97 000.84 000.97 000.93 000.97 000.92 000.86 000.97 Total errors 63.00 38.00 35.00 26.00 27.00 24.00 52.00 117.00 37.00 45.00 67.00 31.00 48.00 49.00 38.00 45.00 46.00 57.00 83.00 32.00 44.00 Incomplete implementations 0.00 17.00 27.00 8.00 10.00 11.00 4.00 33.00 5.00 19.00 40.00 10.00 11.00 9.00 6.00 0.00 20.00 18.00 64.00 20.00 33.00 Runtime errors 7.00 18.00 6.00 17.00 17.00 13.00 23.00 14.00 22.00 26.00 25.00 19.00 36.00 40.00 28.00 35.00 24.00 39.00 19.00 11.00 11.00 Timeouts 56.00 3.00 2.00 1.00 0.00 0.00 25.00 70.00 10.00 0.00 2.00 2.00 1.00 0.00 4.00 10.00 2.00 0.00 0.00 1.00 0.00 Unsatisfiable constraints 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Table 12. Evaluation results of StarCoder2 by API for full completion with unconstrained decoding
+Overall Asana Google Calendar Google Sheets Slack
+Executable implementations (t) 0.99 0.99 1.00 1.00 0.99
+Correct implementations (t) 0.26 0.37 0.46 0.18 0.13
+Correct implementations (e) 0.27 0.37 0.46 0.18 0.13
+Correct URLs (t) 0.60 0.60 0.89 0.47 0.54
+Correct URLs (e) 0.60 0.61 0.89 0.47 0.55
+Illegal URLs (t) 0.22 0.25 0.03 0.18 0.22
+Illegal URLs (e) 0.22 0.25 0.03 0.18 0.23
+Correct methods (t) 0.86 0.90 0.89 0.47 0.86
+Correct methods (e) 0.87 0.90 0.89 0.47 0.87
+Illegal methods (t) 0.11 0.10 0.05 0.29 0.11
+Illegal methods (e) 0.11 0.10 0.05 0.29 0.11
+Correct argument names (t) 0.63 0.74 0.81 0.64 0.49
+Correct argument names (e) 0.63 0.74 0.81 0.64 0.49
+Correct argument values (t) 0.58 0.64 0.79 0.59 0.46
+Correct argument values (e) 0.58 0.64 0.79 0.59 0.47
+Missing arguments (t) 0.37 0.26 0.19 0.36 0.51
+Missing arguments (e) 0.37 0.26 0.19 0.36 0.51
+Unexpected arguments (t) 0.25 0.11 0.18 0.26 0.36
+Unexpected arguments (e) 0.25 0.11 0.18 0.26 0.37
+Mean argument precision (t) 0.68 0.89 0.76 0.64 0.47
+Mean argument precision (e) 0.68 0.89 0.76 0.64 0.47
+Mean argument recall (t) 0.63 0.76 0.75 0.66 0.47
+Mean argument recall (e) 0.63 0.76 0.75 0.66 0.48
+Mean argument Jaccard index (t) 0.56 0.72 0.68 0.52 0.38
+Mean argument Jaccard index (e) 0.56 0.72 0.68 0.52 0.39
+Mean argument value conditional accuracy (t) 0.92 0.87 0.98 0.95 0.96
+Mean argument value conditional accuracy (e) 0.92 0.87 0.98 0.95 0.96
+Total errors 3.00 1.00 0.00 0.00 2.00
+Incomplete implementations 2.00 1.00 0.00 0.00 1.00
+Runtime errors 1.00 0.00 0.00 0.00 1.00
+Table 13. Evaluation results of StarCoder2 by API for argument completion with unconstrained decoding
+Overall Asana Google Calendar Google Sheets Slack
+Executable implementations (t) 0.99 0.99 1.00 1.00 0.99
+Correct implementations (t) 0.25 0.37 0.46 0.24 0.10
+Correct implementations (e) 0.25 0.37 0.46 0.24 0.10
+Illegal implementations (t) 0.35 0.23 0.08 0.24 0.55
+Illegal implementations (e) 0.36 0.23 0.08 0.24 0.55
+Correct argument names (t) 0.64 0.69 0.86 0.67 0.55
+Correct argument names (e) 0.65 0.69 0.86 0.67 0.55
+Correct argument values (t) 0.57 0.59 0.81 0.61 0.49
+Correct argument values (e) 0.58 0.59 0.81 0.61 0.50
+Missing arguments (t) 0.36 0.31 0.14 0.33 0.45
+Missing arguments (e) 0.35 0.31 0.14 0.33 0.45
+Unnecessary arguments (t) 0.07 0.03 0.10 0.06 0.10
+Unnecessary arguments (e) 0.07 0.03 0.10 0.06 0.10
+Illegal arguments (t) 0.14 0.11 0.02 0.06 0.20
+Illegal arguments (e) 0.14 0.11 0.02 0.06 0.20
+Mean argument precision (t) 0.72 0.85 0.82 0.88 0.56
+Mean argument precision (e) 0.72 0.85 0.82 0.88 0.56
+Mean argument recall (t) 0.64 0.72 0.80 0.68 0.53
+Mean argument recall (e) 0.65 0.72 0.80 0.68 0.54
+Mean argument Jaccard index (t) 0.57 0.68 0.74 0.64 0.42
+Mean argument Jaccard index (e) 0.57 0.69 0.74 0.64 0.43
+Mean argument value conditional accuracy (t) 0.90 0.87 0.96 0.90 0.92
+Mean argument value conditional accuracy (e) 0.90 0.87 0.96 0.90 0.92
+Total errors 3.00 1.00 0.00 0.00 2.00
+Incomplete implementations 1.00 0.00 0.00 0.00 1.00
+Runtime errors 2.00 1.00 0.00 0.00 1.00
+Table 14. Evaluation results of StarCoder2 by API for full completion with constrained decoding
+Overall Asana Google Calendar Google Sheets Slack
+Executable implementations (t) 0.90 0.93 1.00 0.94 0.84
+Correct implementations (t) 0.40 0.42 0.51 0.12 0.39
+Correct implementations (e) 0.45 0.45 0.51 0.12 0.46
+Correct URLs (t) 0.61 0.66 0.86 0.47 0.51
+Correct URLs (e) 0.67 0.71 0.86 0.50 0.61
+Illegal URLs (t) 0.00 0.00 0.00 0.00 0.00
+Illegal URLs (e) 0.00 0.00 0.00 0.00 0.00
+Correct methods (t) 0.83 0.86 0.89 0.47 0.83
+Correct methods (e) 0.93 0.92 0.89 0.50 0.99
+Illegal methods (t) 0.00 0.00 0.00 0.00 0.00
+Illegal methods (e) 0.00 0.00 0.00 0.00 0.00
+Correct argument names (t) 0.68 0.71 0.82 0.53 0.64
+Correct argument names (e) 0.77 0.77 0.82 0.56 0.78
+Correct argument values (t) 0.64 0.64 0.81 0.50 0.62
+Correct argument values (e) 0.72 0.70 0.81 0.52 0.76
+Missing arguments (t) 0.32 0.29 0.18 0.47 0.36
+Missing arguments (e) 0.23 0.23 0.18 0.44 0.22
+Unexpected arguments (t) 0.20 0.14 0.27 0.33 0.22
+Unexpected arguments (e) 0.22 0.15 0.27 0.34 0.26
+Mean argument precision (t) 0.78 0.84 0.70 0.58 0.75
+Mean argument precision (e) 0.78 0.84 0.70 0.58 0.75
+Mean argument recall (t) 0.71 0.74 0.75 0.54 0.69
+Mean argument recall (e) 0.79 0.79 0.75 0.58 0.82
+Mean argument Jaccard index (t) 0.63 0.70 0.66 0.43 0.59
+Mean argument Jaccard index (e) 0.70 0.74 0.66 0.46 0.70
+Mean argument value conditional accuracy (t) 0.94 0.89 0.99 0.96 0.98
+Mean argument value conditional accuracy (e) 0.94 0.89 0.99 0.96 0.98
+Total errors 40.00 11.00 0.00 1.00 28.00
+Incomplete implementations 14.00 5.00 0.00 0.00 9.00
+Runtime errors 24.00 4.00 0.00 1.00 19.00
+Timeouts 2.00 2.00 0.00 0.00 0.00
+Unsatisfiable constraints 0.00 0.00 0.00 0.00 0.00
+Table 15. Evaluation results of StarCoder2 by API for argument completion with constrained decoding
+Overall Asana Google Calendar Google Sheets Slack
+Executable implementations (t) 0.93 0.95 0.97 0.94 0.90
+Correct implementations (t) 0.49 0.34 0.49 0.29 0.66
+Correct implementations (e) 0.53 0.36 0.50 0.31 0.73
+Illegal implementations (t) 0.00 0.00 0.00 0.00 0.00
+Illegal implementations (e) 0.00 0.00 0.00 0.00 0.00
+Correct argument names (t) 0.79 0.71 0.84 0.75 0.86
+Correct argument names (e) 0.86 0.75 0.87 0.80 0.98
+Correct argument values (t) 0.75 0.62 0.80 0.69 0.86
+Correct argument values (e) 0.81 0.66 0.83 0.73 0.97
+Missing arguments (t) 0.21 0.29 0.16 0.25 0.14
+Missing arguments (e) 0.14 0.25 0.13 0.20 0.02
+Unnecessary arguments (t) 0.08 0.05 0.14 0.10 0.10
+Unnecessary arguments (e) 0.09 0.05 0.15 0.10 0.11
+Illegal arguments (t) 0.00 0.00 0.00 0.00 0.00
+Illegal arguments (e) 0.00 0.00 0.00 0.00 0.00
+Mean argument precision (t) 0.92 0.96 0.80 0.90 0.91
+Mean argument precision (e) 0.92 0.96 0.80 0.90 0.91
+Mean argument recall (t) 0.81 0.73 0.79 0.75 0.89
+Mean argument recall (e) 0.87 0.77 0.81 0.80 0.98
+Mean argument Jaccard index (t) 0.76 0.72 0.72 0.70 0.81
+Mean argument Jaccard index (e) 0.81 0.75 0.74 0.74 0.90
+Mean argument value conditional accuracy (t) 0.93 0.85 0.97 0.89 0.99
+Mean argument value conditional accuracy (e) 0.93 0.85 0.97 0.89 0.99
+Total errors 27.00 8.00 1.00 1.00 17.00
+Incomplete implementations 10.00 5.00 0.00 1.00 4.00
+Runtime errors 17.00 3.00 1.00 0.00 13.00
+Timeouts 0.00 0.00 0.00 0.00 0.00
+Unsatisfiable constraints 0.00 0.00 0.00 0.00 0.00
+Table 16. Evaluation results of GPT-4o by API for full completion with unconstrained decoding
+Overall Asana Google Calendar Google Sheets Slack
+Executable implementations (t) 1.00 0.99 1.00 1.00 1.00
+Correct implementations (t) 0.60 0.56 0.89 0.59 0.57
+Correct implementations (e) 0.60 0.57 0.89 0.59 0.57
+Correct URLs (t) 0.78 0.75 1.00 0.76 0.75
+Correct URLs (e) 0.78 0.76 1.00 0.76 0.75
+Illegal URLs (t) 0.09 0.12 0.00 0.00 0.10
+Illegal URLs (e) 0.09 0.12 0.00 0.00 0.10
+Correct methods (t) 0.91 0.90 0.89 0.88 0.93
+Correct methods (e) 0.91 0.90 0.89 0.88 0.93
+Illegal methods (t) 0.07 0.10 0.00 0.00 0.07
+Illegal methods (e) 0.07 0.10 0.00 0.00 0.07
+Correct argument names (t) 0.86 0.83 1.00 0.88 0.86
+Correct argument names (e) 0.87 0.84 1.00 0.88 0.86
+Correct argument values (t) 0.83 0.79 1.00 0.81 0.83
+Correct argument values (e) 0.83 0.80 1.00 0.81 0.83
+Missing arguments (t) 0.14 0.17 0.00 0.12 0.14
+Missing arguments (e) 0.13 0.16 0.00 0.12 0.14
+Unexpected arguments (t) 0.11 0.11 0.00 0.06 0.13
+Unexpected arguments (e) 0.11 0.11 0.00 0.06 0.13
+Mean argument precision (t) 0.89 0.90 1.00 0.93 0.85
+Mean argument precision (e) 0.89 0.90 1.00 0.93 0.85
+Mean argument recall (t) 0.87 0.85 1.00 0.89 0.86
+Mean argument recall (e) 0.87 0.86 1.00 0.89 0.86
+Mean argument Jaccard index (t) 0.83 0.81 1.00 0.86 0.81
+Mean argument Jaccard index (e) 0.83 0.81 1.00 0.86 0.81
+Mean argument value conditional accuracy (t) 0.96 0.95 1.00 0.93 0.97
+Mean argument value conditional accuracy (e) 0.96 0.95 1.00 0.93 0.97
+Total errors 1.00 1.00 0.00 0.00 0.00
+Incomplete implementations 0.00 0.00 0.00 0.00 0.00
+Runtime errors 1.00 1.00 0.00 0.00 0.00
+Table 17. Evaluation results of GPT-4o by API for argument completion with unconstrained decoding
+Overall Asana Google Calendar Google Sheets Slack
+Executable implementations (t) 1.00 0.99 1.00 1.00 1.00
+Correct implementations (t) 0.77 0.77 0.86 0.71 0.75
+Correct implementations (e) 0.77 0.77 0.86 0.71 0.75
+Illegal implementations (t) 0.09 0.07 0.14 0.06 0.11
+Illegal implementations (e) 0.09 0.07 0.14 0.06 0.11
+Correct argument names (t) 0.93 0.96 0.93 0.92 0.91
+Correct argument names (e) 0.93 0.97 0.93 0.92 0.91
+Correct argument values (t) 0.90 0.91 0.93 0.89 0.88
+Correct argument values (e) 0.90 0.92 0.93 0.89 0.88
+Missing arguments (t) 0.07 0.04 0.07 0.08 0.09
+Missing arguments (e) 0.07 0.03 0.07 0.08 0.09
+Unnecessary arguments (t) 0.01 0.02 0.00 0.00 0.02
+Unnecessary arguments (e) 0.01 0.02 0.00 0.00 0.02
+Illegal arguments (t) 0.05 0.02 0.07 0.02 0.07
+Illegal arguments (e) 0.05 0.02 0.07 0.02 0.07
+Mean argument precision (t) 0.94 0.97 0.93 0.98 0.90
+Mean argument precision (e) 0.94 0.97 0.93 0.98 0.90
+Mean argument recall (t) 0.93 0.97 0.93 0.93 0.91
+Mean argument recall (e) 0.94 0.97 0.93 0.93 0.91
+Mean argument Jaccard index (t) 0.91 0.95 0.91 0.92 0.88
+Mean argument Jaccard index (e) 0.92 0.95 0.91 0.92 0.88
+Mean argument value conditional accuracy (t) 0.97 0.96 1.00 0.96 0.97
+Mean argument value conditional accuracy (e) 0.97 0.96 1.00 0.96 0.97
+Total errors 1.00 1.00 0.00 0.00 0.00
+Incomplete implementations 0.00 0.00 0.00 0.00 0.00
+Runtime errors 1.00 1.00 0.00 0.00 0.00
+Table 18.
+Relative increase of correct codes from unconstrained decoding (UD) to constrained decoding (CD) for the full and argument completion, relative to all samples (t) and relative to executable codes (e).
+Models with a 0% unconstrained baseline performance are excluded from the calculation of the average.
+(a) Full completion (t)
+UD CD Gain
+CodeT5+ (16B) 0.06 0.13 +126%
+StarCoder (15.5B) 0.12 0.15 +24%
+StarCoder2 (3B) 0.11 0.18 +62%
+StarCoder2 (7B) 0.10 0.18 +71%
+StarCoder2 (15B) 0.26 0.40 +52%
+DeepSeek-Coder (1.3B) 0.08 0.18 +135%
+DeepSeek-Coder (6.7B) 0.17 0.29 +65%
+DeepSeek-Coder (7B) 0.04 0.08 +136%
+DeepSeek-Coder (33B) 0.17 0.39 +129%
+DeepSeek-Coder-V2 (16B) 0.21 0.27 +31%
+Qwen2.5-Coder (0.5B) 0.00 0.04 +inf%
+Qwen2.5-Coder (1.5B) 0.00 0.02 +inf%
+Qwen2.5-Coder (3B) 0.00 0.07 +inf%
+Qwen2.5-Coder (7B) 0.00 0.30 +inf%
+Qwen2.5-Coder (14B) 0.00 0.07 +inf%
+Qwen2.5-Coder (32B) 0.00 0.36 +inf%
+Llama 3.1 (8B) 0.00 0.19 +inf%
+Llama 3.1 (70B) 0.00 0.31 +inf%
+Code Llama (7B) 0.12 0.23 +85%
+Code Llama (13B) 0.08 0.25 +200%
+Code Llama (70B) 0.30 0.46 +52%
+Average 0.09 0.22 +90%
+(b) Full completion (e)
+UD CD Gain
+CodeT5+ (16B) 0.06 0.16 +158%
+StarCoder (15.5B) 0.13 0.17 +33%
+StarCoder2 (3B) 0.12 0.23 +95%
+StarCoder2 (7B) 0.11 0.21 +87%
+StarCoder2 (15B) 0.27 0.45 +68%
+DeepSeek-Coder (1.3B) 0.08 0.21 +160%
+DeepSeek-Coder (6.7B) 0.18 0.35 +93%
+DeepSeek-Coder (7B) 0.06 0.12 +94%
+DeepSeek-Coder (33B) 0.17 0.44 +152%
+DeepSeek-Coder-V2 (16B) 0.21 0.32 +50%
+Qwen2.5-Coder (0.5B) 0.00 0.06 +inf%
+Qwen2.5-Coder (1.5B) 0.00 0.02 +inf%
+Qwen2.5-Coder (3B) 0.00 0.09 +inf%
+Qwen2.5-Coder (7B) 0.00 0.36 +inf%
+Qwen2.5-Coder (14B) 0.00 0.08 +inf%
+Qwen2.5-Coder (32B) 0.00 0.46 +inf%
+Llama 3.1 (8B) 0.00 0.24 +inf%
+Llama 3.1 (70B) 0.00 0.38 +inf%
+Code Llama (7B) 0.12 0.31 +148%
+Code Llama (13B) 0.09 0.28 +204%
+Code Llama (70B) 0.31 0.54 +78%
+Average 0.09 0.26 +109%
+(c) Argument completion (t)
+UD CD Gain
+CodeT5+ (16B) 0.12 0.35 +180%
+StarCoder (15.5B) 0.27 0.31 +17%
+StarCoder2 (3B) 0.21 0.37 +77%
+StarCoder2 (7B) 0.19 0.30 +57%
+StarCoder2 (15B) 0.25 0.49 +96%
+DeepSeek-Coder (1.3B) 0.21 0.39 +84%
+DeepSeek-Coder (6.7B) 0.26 0.47 +80%
+DeepSeek-Coder (7B) 0.08 0.23 +181%
+DeepSeek-Coder (33B) 0.25 0.58 +136%
+DeepSeek-Coder-V2 (16B) 0.35 0.48 +36%
+Qwen2.5-Coder (0.5B) 0.08 0.14 +77%
+Qwen2.5-Coder (1.5B) 0.03 0.11 +231%
+Qwen2.5-Coder (3B) 0.16 0.27 +69%
+Qwen2.5-Coder (7B) 0.28 0.51 +81%
+Qwen2.5-Coder (14B) 0.37 0.46 +24%
+Qwen2.5-Coder (32B) 0.38 0.65 +71%
+Llama 3.1 (8B) 0.05 0.47 +879%
+Llama 3.1 (70B) 0.29 0.51 +78%
+Code Llama (7B) 0.23 0.38 +62%
+Code Llama (13B) 0.12 0.44 +258%
+Code Llama (70B) 0.40 0.63 +55%
+Average 0.22 0.41 +135%
+(d) Argument completion (e)
+UD CD Gain
+CodeT5+ (16B) 0.13 0.41 +213%
+StarCoder (15.5B) 0.27 0.34 +26%
+StarCoder2 (3B) 0.22 0.41 +90%
+StarCoder2 (7B) 0.19 0.32 +66%
+StarCoder2 (15B) 0.25 0.53 +109%
+DeepSeek-Coder (1.3B) 0.22 0.41 +91%
+DeepSeek-Coder (6.7B) 0.27 0.55 +105%
+DeepSeek-Coder (7B) 0.08 0.32 +295%
+DeepSeek-Coder (33B) 0.25 0.64 +159%
+DeepSeek-Coder-V2 (16B) 0.36 0.55 +52%
+Qwen2.5-Coder (0.5B) 0.09 0.17 +89%
+Qwen2.5-Coder (1.5B) 0.03 0.12 +242%
+Qwen2.5-Coder (3B) 0.18 0.31 +77%
+Qwen2.5-Coder (7B) 0.28 0.58 +106%
+Qwen2.5-Coder (14B) 0.40 0.50 +28%
+Qwen2.5-Coder (32B) 0.38 0.73 +92%
+Llama 3.1 (8B) 0.05 0.53 +983%
+Llama 3.1 (70B) 0.29 0.59 +105%
+Code Llama (7B) 0.23 0.48 +103%
+Code Llama (13B) 0.13 0.47 +271%
+Code Llama (70B) 0.40 0.70 +74%
+Average 0.22 0.46 +161%